### v1

# An Ran Chen

Abstract
Bug fixing is a crucial task in software maintenance to hold user trust. Although various automated fault localization techniques exist, they often require specific conditions to be effective. For example, Spectrum-Based Fault Localization (SBFL) techniques need at least one failing test to identify bugs, which may not always be available. Bug reports, particularly those with stack traces, provide detailed information on system execution failures and are invaluable for developers. This study focuses on utilizing stack traces from crash reports as fault-triggering tests for SBFL. Our findings indicate that only 3.33% of bugs have fault-triggering tests, limiting traditional SBFL efficiency. However, 98.3% of bugfix intentions align directly with exceptions in stack traces, and 78.3% of buggy methods are reachable within an average of 0.34 method calls, proving stack traces as a reliable source for locating bugs. We introduce a new approach, SBEST, that integrates stack trace data with test coverage to enhance fault localization. Our approach shows a significant improvement, increasing Mean Average Precision (MAP) by 32.22% and Mean Reciprocal Rank (MRR) by 17.43% over traditional stack trace ranking methods.
Keywords:
bug report, stack trace, empirical study, fault localization, software maintenance.
1 Introduction
Locating and fixing software bugs is an essential yet frequently expensive and time-consuming task. Bug-fixing activities not only delay the project’s progress but also increase the software’s overall cost, making them a critical part of software maintenance. Considering the complexity and rapid evolution of modern software systems, the ability to efficiently and accurately locate bugs is more important than ever. To address this challenge, previous research has proposed numerous different automated fault localization techniques (Abreu et al., 2007; Jones and Harrold, 2005; Naish et al., 2011; Abreu et al., 2009; Wong et al., 2013; Papadakis and Le Traon, 2015; Hong et al., 2017; Papadakis and Le Traon, 2012; Wang et al., 2022; Cui et al., 2020), among which Spectrum-Based Fault Localization (SBFL) approaches are prominent due to its high accuracy (de Souza et al., 2016). SBFL techniques leverage information derived from test executions — specifically, the spectrum of code elements executed during passing and failing tests — to calculate the probability of every code element being faulty (i.e. suspiciousness score).
The effectiveness of SBFL highly depends on the availability of failing test cases that uncover the bug (i.e., fault-triggering tests). However, studies indicate that in real-world scenarios, especially in continuous deployment and integration environments, failing tests may not always be present or identified when bugs are reported (Chen et al., 2023; Kabadi et al., 2023). This absence of fault-triggering tests significantly impacts the applicability of SBFL, as it depends on it to distinguish between faulty and non-faulty code. In the absence of failing tests, alternative approaches must be considered. Exception stack traces, commonly included in crash reports, provide dynamic execution information that can be extremely valuable for debugging. Studies show that developers usually rely on them when investigating bugs (Schroter et al., 2010; Zimmermann et al., 2010). In addition, previous research (Chen et al., 2021a) shows that stack traces provide valuable information on the cause of the bugs and their location in the source code. These stack traces, which outline the sequence of method calls leading up to an exception, offer implicit clues about where a bug may reside, providing a fallback localization method in environments where traditional test-based approaches falter.
In this paper, we address the challenges of fault localization in scenarios where traditional fault-triggering test data is scarce. We conduct our study on the Defects4J 2.0 dataset (Just et al., 2014b), a widely recognized resource for research in fault localization and program repair. However, prior studies (Kabadi et al., 2023; Chen et al., 2023) have reported that some tests in Defects4J were added post hoc, following the reporting of bugs, which could bias the bug localization process. To avoid this issue and better simulate the state of the codebase at the time developers first address a bug, we utilize the commit data immediately preceding the bug report rather than relying directly on the buggy commits provided by Defects4J. Our initial analysis of the bugs reveals a critical limitation: only 3.33% of crash reports, which are bug reports that contain stack traces, possess fault-triggering tests. Even for bug reports without stack traces, only 10.19% of the bugs have fault-triggering tests. This low incidence of failing tests challenges the efficacy of Spectrum-Based Fault Localization (SBFL) methods, which rely heavily on test outcomes to identify bug locations. To overcome this limitation, we propose leveraging the stack traces—rich in contextual and execution information about the bug’s root causes—as a substitute for the failing tests. We introduce a novel approach, Spectrum-Based Localization Enhanced by Stack Traces (SBEST), which integrates stack trace analysis within the SBFL framework. Our method significantly improves the accuracy of fault localization by utilizing the comprehensive code execution information allied with the exception stack traces. This integration enables a more detailed understanding of system behaviour at the moment of failure, thereby addressing the gaps left by the unavailability of failing tests.
In particular, we study and answer three research questions (RQs):
  • RQ1: Are the test failures related to the bug in crash reports? We study the presence of fault-triggering tests in crash reports (i.e., bug reports that contain stack traces) and their impact on the SBFL results. We find that only 3.33% of the crash report bugs contain fault-triggering tests, which profoundly degrades the SBFL efficiency in locating these bugs. The limited execution data underscores the importance of incorporating stack traces into the SBFL process, enhancing the potential to accurately identify bug locations in the absence of traditional test failures.
  • RQ2: What is the relationship between stack traces and buggy location? Given that the stack traces are the only execution information available in most of the studied bug reports, we aim to study how they relate to the buggy location and how they can be used to locate these bugs. To do so, we inspect the type of modification performed in the bugfix commit and the distance between the stack trace and the buggy methods. We find that in 98.3% of the studied bugs, the bugfix intention is directly correlated with the exception in the stack trace (e.g., adding code to handle the exception). In addition, 78.3% of the buggy methods are directly reachable from the stack traces, with an average distance of 0.34 method calls. This indicates that leveraging the stack traces to reconstruct the execution path at the time of the exception can be an effective method for pinpointing the bugs.
  • RQ3: Can we utilize the stack traces to help detect the buggy locations? Given the previous findings, we propose a new approach, SBEST, that integrates the stack trace information into Spectrum-Based Fault Localization. We compare our approach with two baselines: the Ochiai results and the position of the methods in the stack trace. Ochiai proves to be the least effective technique by a significant margin. On the other hand, we find that the stack traces alone can locate 34 out of the 60 bugs in the Top-5. This reinforces that the stack traces are a vital source of information about the bug location and should be prioritized in FL. In addition, SBEST achieves an improvement of 32.22% on MAP and 17.43% on MRR compared to the Stack Trace ranking. The enhancement of metrics suggests that the incorporation of coverage information effectively improved the ranking of stack traces, thereby providing the tool with a more comprehensive view of system execution at the time of failure.
In summary, this paper presents a comprehensive study that evaluates the effectiveness of combining stack traces with test coverage data. Our findings indicate that by incorporating stack traces into the SBFL model, we can significantly improve bug localization, thereby addressing one of the key challenges in automated debugging processes. We make the data publically available online for replication and encourage future research in this area (sbe, 2024).
Paper Organization. Section 2 details the background and the motivation behind this study. Section 3 surveys the related work. Section 4 explains the experimental setup. Section 5 presents the results. Section 6 discusses the implications of our findings. Section 7 presents the threats to validity. Section 8 concludes the paper and discusses future work.
2 Background and Motivation
Bug Reports contain important information for developers to fix the issues that users or other developers encounter in the software (Chen et al., 2021a). A bug report typically includes a title, a description, attached files, and comments. For example, Figure 1 shows a bug report from the Defects4j 2.0 dataset. The title offers a concise one-line description of the problem. The description provides developers with the necessary details to investigate the bug, such as the stack traces illustrating the system execution details. When additional information is required, users and developers use the comments and attached files sections to further communicate on the bug. In Figure 1 bug COMPRESS-181, a problem occurs when the system reads TAR files with symlinks. In particular, the description provides developers with a stack trace that illustrates an IOException is being thrown. The description also provides the affected nightly builds, which show no test failures.
Figure 1: An example of a bug report (COMPRESS-181). The bug report addresses a problem in reading TAR files that contain symlinks, which cause an IOException to be thrown. The report contains the stack trace of the error (gray box), and an example of TAR file (simple-aix-native-tar.tar) that triggers the bug is attached. When this bug report was created, there were no failing tests.
To assist developers in debugging such bugs, prior studies have proposed Spectrum-Based Fault Localization (SBFL) techniques (Chen et al., 2022; Lou et al., 2021; Baudry et al., 2006). SBFL relies on the information from failing tests to identify potentially faulty code. Hence, SBFL assumes that test failures that are relevant to the bug exist. SBFL works by comparing the code execution of passing and failing tests to identify their differences. The potential problematic code is then identified based on these differences. SBFL is widely studied in prior studies for its simplicity and efficiency (Santelices et al., 2009; Pearson et al., 2016). However, when there is no failing test, SBFL can not distinguish buggy code elements. For instance, in the case of bug COMPRESS-181 (Figure 1), in which there are no failing tests, SBFL can not help prioritize the faulty locations. This is because, in the absence of failing tests, there is no coverage spectrum between passing and failing executions. Thus, it becomes impossible to prioritize any code element for investigation, as SBFL relies on failing executions as a guide to identify where faults might be present. In this case, all the code locations receive the exact same suspiciousness score of zero, and the bug cannot be located.
In this paper, we investigate in detail the characteristics of Defects4j 2.0 crash reports (i.e., bug reports that contain stack traces). We dissect information such as the test failures, coverage, and stack traces by conducting a quantitative and qualitative analysis of the data and propose a new approach to address the limitations found in SBFL in such scenarios.
3 Related Work
3.1 Fault Localization
Automated fault localization methods are a recurrent theme in the literature, being Spectrum-Based Fault Localization (SBFL) and Mutation Based Fault Localization (MBFL) two of the most famous representatives. SBFL is a very consolidated class of Fault Localization techniques. These approaches use the test information to localize the bugs, and they require at least one failing test to work. Ochiai (Abreu et al., 2007) is one of the most famous SBFL techniques. Other representatives of this class are Tarantula (Jones and Harrold, 2005), Op2 (Naish et al., 2011), BARINEL (Abreu et al., 2009) and DStar (Wong et al., 2013): similar approaches based on different formulas to compute the suspiciousness score. Given that SBFL techniques are light weighted and fast, they are also applied to other kinds of techniques such as interactive fault-localization (Gong et al., 2012) and program repair (Ye et al., 2022; Le Goues et al., 2012).
MBFL techniques (Papadakis and Le Traon, 2015; Moon et al., 2014; Hong et al., 2017; Papadakis and Le Traon, 2012; Wang et al., 2022) are based on mutation analysis. For each part of the code covered by failed test cases, a set of mutants is created, and the test suite is executed. Based on the test results, the suspiciousness scores are calculated, and the most suspicious locations are ranked. Despite presenting promising results, this kind of technique is very costly and time-consuming since it requires multiple executions of the test suite. Strategies for reducing this overhead and improving the general accuracy have been studied, such as ways prioritizing the mutants (Liu et al., 2018), prioritizing the tests executed for each mutant (de Oliveira et al., 2018) and a mix of spectrum and mutant-based approaches (Cui et al., 2020). One common point between SBFL and MBFL approaches is that both types of techniques require failing tests to work. However, this might be not always the case in real life.
Most of the mentioned studies use Defects4j to evaluate the results. Defects4j is one of the most famous benchmarks for test-based fault localization research. Nonetheless, recent studies (Kabadi et al., 2023; Chen et al., 2023) show that often in Defects4j, tests that were created after the bug-report creation are added to the buggy patch and thus contain developer knowledge about the bug investigation process. These tests were artificially appended to the buggy commit in Defects4j to simulate the scenario of a bug with test failures. This fact indicates that many fault localization techniques that were evaluated in Defects4j might have very different results when applied to real-world scenarios.
3.2 Bug Reports and Stack Traces
The bug report quality is a common topic in previous research. Knowing which parts of the report are more important in the bug-fixing process is very important to guide new studies on fault localization. Our focus on the stack trace information is supported by some previous results.
In the work by Bettenburg et al. (2008), the authors conduct a survey to understand what makes a good bug report from the point of view of the developers. They found that information about the steps to reproduce the bug, stack traces, and test cases are considered very helpful. In the work by Schroter et al. (2010), the authors analyze the intersection between the methods in the stack trace and the methods modified in the bugfix commit. They found out that more than 47% of the stack traces extracted from bug reports contained at least one buggy method.
An empirical study from Chen et al. (2021a) analyzes the logs’ (stack traces and log snippets) importance in crash reports. By extensively analyzing 1,561 bug reports with logs, they concluded that logs often provide helpful information about where the bug is located. They found out that in 73% of the bug reports, there is an overlap between the classes in the logs and the fixed classes.
4 Data Collection and Case Study Setup
In this section, we start by describing the data collection process and presenting our dataset. After this, we detail the evaluation metrics.
4.1 Data Collection
4.1.1 Collecting Bug Reports with Stack Traces
In this paper, we study crash reports (i.e., bug reports that contain stack traces) in 15 projects from Defects4j version 2.0.0. Defects4j (Just et al., 2014b) is a benchmark and framework of real bugs used in many previous Software Engineering studies related to bug repair and fault localization (Lutellier et al., 2020; Chen et al., 2021b; Li et al., 2021; Ye et al., 2022). The bugs in Defects4j comprise a wide variety of systems with different characteristics, with sizes ranging from 4K to 90K lines of code (LOC) and the number of tests varying from 54 to 7,911. Although Defects4J provides some basic information about the bugs (e.g., bug report ID and the commit hash of the fixes), it does not contain the stack traces provided in the bug reports. Hence, we start by extracting the bug report URL for all the bugs in the Defects4J repository. To collect the bug reports’ textual information (title, description, and comments) and their creation date, we implement a crawler to retrieve information from the project management tool of each system (Jira, GitHub Issues, or Source Forge).
Figure 2: An example of the structure of stack trace entries in Java.
In total, we collected 803 bugs for which the corresponding bug reports were available. For each bug, we then combine the bug report information with the information available in Defects4j, including the bugfix commit hash and the list of fault-triggering tests. Fault-triggering tests are defined as tests that fail in the buggy commit and pass in the bugfix commit. They are tests that cover the buggy code and are essential for pinpointing the buggy location when using SBFL approaches. As shown in Figure 2, stack traces have a pre-defined pattern. Therefore, we implement a regex-based parser to identify stack traces in bug reports. Our parser identifies the stack traces and extracts the file name, method name, and line number from each stack trace entry. Among all the 803 bugs we collected from Defects4J, we identified 89 bugs that have stack traces in the bug report.
Table 1: An overview of our studied systems from Defects4J v2.0.0. #Total Bugs, #Bugs with Stack Traces, #Studied Bugs, LOC, and #Tests show the total number of bugs in Defects4j, the total number of bugs with stack traces, the number of bugs studied, lines of code, and tests in each system, respectively.
System #Total #Bugs with #Studied LOC #Tests
Bugs Stack Traces Bugs
Cli 39 3 2 4K 94
Closure 174 8 8 90K 7,911
Codec 18 1 1 7K 206
Collections 4 1 1 65K 1,286
Compress 47 11 11 9K 73
Csv 16 1 2 2K 54
Gson 18 3 3 14K 720
JacksonCore 26 4 4 22K 206
JacksonDatabind 112 27 10 4K 1,098
Jsoup 93 8 8 8K 139
JxPath 22 1 1 25K 308
Lang 64 4 4 22K 2,291
Math 106 6 3 85K 4,378
Mockito 38 8 2 11K 1,379
Time 26 2 2 28K 4,041
Total 803 89 60 380K 24,302
4.1.2 Collecting Test Coverage
To obtain the test information for fault localization, we need to collect the test execution results, detailed test coverage, and the bug-fix patch. However, there are some limitations in Defects4j. Recent studies (Kabadi et al., 2023; Chen et al., 2023) found that many bugs in Defects4J v1.0 contain tests from the “future” (i.e., added by developers after the bugs were fixed) in their designated buggy commits. This is problematic because the test coverage may contain developers’ knowledge of the bug, which can cause noise and bias in the result of the downstream research. Because of this, we do not directly utilize Defects4J’s data in our study. Instead, since we have the creation date of each bug report, we extract the commit right before the bug report creation from the project’s repository. This commit represents the code that was available to developers when they started to address the bug report. We refer to this commit as a bug report commit to differentiate it from the buggy commit provided by Defects4J.
Due to the above-mentioned issues in Defects4J, our next step is to collect the test results, and detailed test coverage for the bug report commits. To do that, we utilized GZoltar (Campos et al., 2012). GZoltar is a debugging, fault localization, and coverage extraction tool for Java applications that was often applied to analyze Defects4J bugs (Silva et al., 2021; Zhang et al., 2021; Pearson et al., 2017; Küçük et al., 2019). For each crash report, we check out the bug report commit, compile the system, and execute the tests. We utilize the GZoltar CLI tool to implement a script to extract the test coverage and test execution results for the bugs from our study. Since there are many ancient commits, we encountered many challenges in compiling the system and running the tests. For example, since some commits in a project can use different versions of JVM, we had to implement a tool to automatically switch JVM versions when analyzing the commits. There are also many dependency-related issues, where a project uses different versions of a library across commits, but upgrading/downgrading the library version can cause dependency conflicts that lead to compilation errors. In total, we spent over 100 hours compiling and collecting the data. Despite our best efforts, we excluded some bugs where we were not able to execute and collect the test coverage information.
Table 1 provides detailed information on the bugs studied in this work, i.e., bugs from Defect4J 2.0 with valid bug reports. In total, we include 803 bugs from 15 open-source projects. In this work, we target user-reported bugs with stack traces reported, i.e., crash-reporting bugs. Among the 15 projects, there are 89 crash-reporting bugs (i.e. bugs with stack traces in their bug reports), which represents 11.08% of the total set. Of these 89 bugs, we were able to compile and collect the test execution information for 63 bugs. Finally, we excluded three bugs from our study whose root cause is not in a method, as this work focuses on method-level bug localization. The absence of buggy methods can happen in two different situations: (1) The bugfix involves altering lines outside the methods (for example, changing the value of a global variable); or (2) The bugfix includes the creation of new methods, but no existing method is updated. In total, we study 60 bugs. Despite the low percentage of crash reports found in Defects4j, they match with the percentage of bugs with stack traces in other projects (Chen et al., 2021a).
4.2 Evaluation Metrics
In this paper, we conduct a series of experiments on locating bugs using stack traces. In order to evaluate and compare the results, we utilize a set of well-consolidated metrics in the fault localization field. Previous studies suggest that performing fault-localization at the file level lacks precision (Kochhar et al., 2016) while opting for the statement-level granularity can cause the approach to miss important code context (Parnin and Orso, 2011). Hence, in this paper, we conduct our analysis at the method level.
Mean Average Precision (MAP). MAP is a metric that considers the rank of all the buggy methods in a given project. It is calculated by taking the mean of the Average Precision (AP) across all these bugs, where the following formula defines the AP, and $m$ is the set of buggy files from a given bug report.
$AP=\frac{\sum_{i=1}^{m}{i}/{Pos(i)}}{m}$
Mean Reciprocal Rank (MRR). MRR considers the position where the first buggy file was ranked and takes the mean from all the bug reports. It is computed by the following formula, where K is the set of bug reports and $rank_{i}$ is the position of the first buggy method in the rank.
$MRR=\frac{1}{K}\sum_{i=1}^{K}\frac{1}{rank_{i}}$
Top K. The Top K metric represents the number of bugs in which at least one buggy method is located between the first K best-ranked methods in the approach. According to a previous study (Parnin and Orso, 2011), developers only check a limited amount of suspicious locations. Due to this, we use 1, 3, and 5 as the values for K.
5 Case Study Results
In this section, we discuss the results of our research questions (RQs). For each RQ, we present the motivation, our approach and the results.
RQ1 - Are the test failures related to the bug in crash reports?
Motivation. Prior research (Abreu et al., 2007; Jones and Harrold, 2005; Naish et al., 2011) uses SBFL to assist in debugging, which relies on failing test coverage. The assumption behind SBFL is that the system has at least one failing test covering the bug (i.e. fault-triggering failing test). Intuitively, if fault-triggering tests fail on the bug report commits (the nearest commit when a bug report was reported), we can leverage SBFL techniques to pinpoint the buggy locations. However, it is not clear if fault-triggering tests are available and failing when a bug report is created (Just et al., 2014a; Haben et al., 2023). Therefore, in this RQ, we study whether there are fault-triggering tests (and if these tests fail) at the moment of the bug report creation and their impact on the SBFL’s efficiency.
Approach. To understand whether the test failures are related to the bugs in crash reports, we analyze the number of crash reports that contain failing fault-triggering tests. As discussed in Section 4.1.2, for each bug report, we perform our study on the bug report commits to avoid biases of ”future” tests. In particular, we execute the fault-triggering tests provided by Defects4J and collect the test results with GZoltar. We also investigate whether fault-triggering tests are related to other types of bug reports (without stack traces) and compare the test results with crash reports. A prior study (Chen et al., 2023) identified a set of bugs without stack traces from Defects4J. Therefore, we used their dataset and identified a total of 157 bugs without stack tracks. This data will allow us to draw a comparison between crash report bugs and bugs without stack traces.
To study the impact of missing failing fault-triggering tests on SBFL techniques, we select Ochiai (Abreu et al., 2007) as the baseline since it is widely used and has been shown to perform well on real faults (Pearson et al., 2017; Le et al., 2013; Abreu et al., 2006). We use the detailed coverage and the test results obtained via the Gzoltar execution to apply the Ochiai formula at the method level. The formula calculates the suspiciousness of each code statement, which allows us to rank the methods based on how probable they contain the bug. The SBFL suspiciousness formulas vary for each technique, but all are based on the idea that the code that is more covered by failing tests and less covered by passing tests is more likely to be buggy. The Ochiai formula assigns a suspiciousness score between 0 (not suspect) and 1 (highly suspect) for each piece of code element (e.g., class, method, or statement). If applied at a method level, we calculate the suspicious score of a given method j as:
$s_{o}=\frac{n_{11}(j)}{\sqrt{(n_{11}(j)+n_{01}(j))*(n_{11}(j)+n_{10}(j))}}$ (1)
Table 2: Description of the terms in the Ochiai formula. Covered indicates if the component (in our case, the method) was executed or not during the testing (i.e., is covered or not by the test). Test indicates if the test case failed or passed during its execution.
n Covered Test
$n_{00}$ no passed
$n_{10}$ yes passed
$n_{01}$ no failed
$n_{11}$ yes failed
in which the terms are as defined in Table  2. Each term n in the formula (e.g., $n_{00}$) corresponds to the number of tests that obey specific criteria for the method j for which the suspiciousness score is being calculated. The first criterion determines the method coverage, whether the method is covered by a test ($n_{1x}$) or not ($n_{0x}$). The second criterion indicates the execution status, passing ($n_{x0}$) or failing ( $n_{x1}$). So, if a test covers the method j and its execution fails, for instance, this test will be computed under $n_{11}$. As a calculation example, suppose that we have the following test case results for a given method M1:
  • 6 failing test cases that cover M1 ($n_{11}$)
  • 2 failing test cases that do not cover M1 ($n_{10}$)
  • 10 passing test cases that cover M1 ($n_{01}$)
By applying Equation 1, we get that the Ochiai score for method M1 is approximately 0.530.
Results. Only 3.33% of the crash report bugs contain fault-triggering tests (failing test cases that uncover the bug). Table 3 shows the test execution and fault-triggering test results for the crash report bugs. When comparing with the bugs without stack traces, we obtain the following numbers:
  • Crash Reports: 2 out of 60 bugs have fault-triggering tests (3.33%)
  • Bugs without Stack Traces: 16 out of 157 bugs have fault-triggering tests (10.19%)
Table 3: Test execution results for the crash report bugs. #Bugs, #Tests and #Bugs with Fault-triggering Tests show the total number of studied bugs, the total amount of tests and the total number of bugs with fault-triggering tests; while Failing Tests (Avg.), and Fault-triggering Tests (Avg.) represent the average number of failing tests and fault-triggering tests, respectively, calculated among the bugs.
System (#Bugs) #Tests Failing Tests Fault-triggering #Bugs with Fault-
(Avg.) Tests (Avg.) Triggering Tests
Cli (2) 94 2 0 0
Closure (8) 7,911 2.125 0 0
Codec (1) 206 0 0 0
Compress (11) 73 1 0 0
Csv (1) 54 0 0 0
Gson (3) 720 0 0 0
JacksonCore (4) 206 0.5 0.25 1
JacksonDatabind (10) 1,098 26.9 0.1 1
Jsoup (8) 139 0.125 0 0
JxPath (1) 308 0 0 0
Lang (4) 2,291 8.5 0 0
Math (3) 4,378 0.667 0 0
Mockito (2) 1,379 21.5 0 0
Time (2) 4,041 20.5 0 0
Total (60) 24,302 5.987 0.025 2
Based on the findings, we observed that the majority of the crash report bugs (96.67%) either lack fault-triggering tests or do not trigger them (i.e., fault-triggering tests did not fail). In addition, we can see that the percentage of fault-triggering tests in crash report bugs is significantly less when compared to the set of bugs without stack traces (3.33% versus 10.19%). The reason may be that such exception-related issues are more often triggered during production and not during testing. Another reason may be that developers are more likely to handle exceptions for debugging purposes rather than fault prevention (Shah et al. (2008)). This shift in the use of exception handling as debugging aids increases the likelihood of exception-related issues occurring in production. In such cases, applying traditional SBFL techniques would not be ideal, given the incapacity of the tool to differentiate buggy statements when there are no test failures covering them. Due to the absence of failing fault-triggering tests, the runtime information included in the bug report (i.e., stack traces) serves as the final resource for assistance. In short, studying how to leverage stack traces for fault localization is an important supplement to traditional SBFL techniques that use failing fault-triggering tests.
SBFL techniques may not be effective in addressing crash reports due to the limited presence of fault-triggering failing tests (3.33% of all bug reports).
To study how the absence of fault-triggering tests impacts SBFL, we apply Ochiai to our studied bugs with stack traces. To evaluate the approach, we utilize the metrics presented in Section 4.2, specifically Top-1, Top-3, Top-5, MAP and MRR. We exclude from the Ochiai results the bugs without failing tests, given that the presence of failing tests is necessary for the computation of the suspicioness score. Out of 60 bugs, 23 did not have any failing tests. Note that even if there is a failing test, the test may not be related to the bug that we are interested in (i.e., it is not a fault-triggering test).
On the bug report commit, SBFL performed poorly in all the projects, only being able to locate 2/60 bugs among the Top-5 methods. Table 4 shows the SBFL results applied to the bug report commit for the crash report bugs that contain failing tests. Ochiai could not locate any bug in Top-1 and located only 1 in Top-3 and 2 in Top-5. The highest MAP and MRR values across projects are 0.0693 and 0.0836, respectively, both in the JacksonCore project. Considering that the values of MAP and MRR range between 0 and 1, the obtained values are extremely low. In contrast, a prior study by Chen et al. (2022) that evaluates Ochiai on the original Defects4J benchmark achieved an average MAP of 0.30 and an average MRR of 0.42, which is significantly higher than our results obtained when there is no failing fault-triggering tests.
The MAP and MRR values assess the effectiveness of fault localization techniques in returning relevant results in the top ranking. These metrics are important indicators that demonstrate the usefulness of fault localization techniques. According to a prior survey by Kochhar et al. (2016), 80% of developers consider a fault localization technique successful if it can localize bugs in the top 5 positions. Hence, our finding shows the inefficacy and limitation of SBFL techniques due to the lack of failing tests related to the bug reports.
Table 4: Ochiai results for all the studied systems.#Bugs represents the total number of studied bugs, while #Bugs with FT is the number of bugs in which there is at least one Failing Test.
System (#Bugs — #Bugs with FT) Top-1 Top-3 Top-5 MAP MRR
Cli (2 — 2) 0 0 0 0.0045 0.0045
Closure (8 — 8) 0 0 0 0.0007 0.0012
Codec (1 — 0) - - - - -
Collections (0 — 0) - - - - -
Compress (11 — 6) 0 0 1 0.0079 0.023
Csv (1 — 0) - - - - -
Gson (3 — 0) - - - - -
JacksonCore (4 — 2) 0 1 1 0.0693 0.0836
JacksonDatabind (10 — 8) 0 0 0 0.0005 0.0005
Jsoup (8 — 1) 0 0 0 0 0
JxPath (1 — 0) - - - - -
Lang (4 — 4) 0 0 0 0.0008 0.0008
Math (3 — 2) 0 0 0 0.0002 0.0001
Mockito (2 — 2) 0 0 0 0.0060 0.0281
Time (2 — 2) 0 0 0 0.0003 0.0003
Due to a lack of failing fault-triggering tests, traditional SBFL techniques have inferior localization results.
RQ2 - What is the relationship between stack traces and buggy location?
Motivation. In the previous RQ, we observed that most crash report bugs do not have fault-triggering tests. Because of that, the stack traces stand out as the best source of execution information. In this RQ, we want to study how stack traces are related to the buggy location and how they can be used to locate these bugs. To do so, we look at two things: (i) what is the type of modification performed in the bug fix to resolve the exception, and (ii) how far away the buggy methods (i.e., methods updated in the bug fix) are from the methods in the stack traces.
Approach. To study the type of modification performed in the bugfix to handle the exception, we manually examine the bug-fix patch of all the studied bugs. Based on the modification performed, we classify the bug-fix intention type into four categories:
  1. 1.
    Exception Prevention: This category includes code modifications to prevent a specific exception’s recurrence. For example, the bugfix patch from Cli-5 introduced a new conditional structure to avoid the occurrence of the reported NullPointerException (Figure 3).
  2. 2.
    Exception Conversion: In this type of bugfix, the exception is converted into a warning or error message. For instance, to fix the bug Closure-152 (Figure 4), the developers handled the ClassCastException being thrown and created a warning detailed message.
  3. 3.
    Exception Wrapping: This category represents the cases in which the developers fix the bug by wrapping the exception into another exception type. One example is the bug COMPRESS-12, in which the exception IllegalArgumentException was wrapped into the exception IOException (Figure 5)
  4. 4.
    Keep Throwing: Finally, this category includes the bugs in which no exception handling-related code is found in the bugfix.
Figure 3: Extract from the bugfix patch for Cli-5, classified in the Exception Prevention category.
Figure 4: Extract from the bugfix patch for Closure-152, classified in the Exception Conversion category.
Figure 5: Extract from the bugfix patch for COMPRESS-12, classified in the Exception Wrapping category.
To study the distance between the stack trace and the buggy methods, we utilize the stack trace content and the code in the bug report commit to create a call graph for each bug. We then measure the minimum distance between a method in the stack trace and one of the buggy methods, if reachable. For example, the bug Math-79 in Figure 6 has a 3-call distance between the stack trace and the buggy method. The distance is zero for a bug if the buggy method is recorded in the stack trace.
Figure 6: Call Graph for the bug Math-79, which has a 3-call distance between the stack trace and the buggy method.
Results. In 83% of the cases, the bugfix intention is to prevent the exception from happening again. Figure 7 shows the bugfix intention type distribution for studied bugs. In the majority of the fixes (83.3%), the intention behind the fix is Exception Prevention (i.e., to prevent the exception from happening again). This means that the stack traces are directly related to the root cause of the bug, which suggests that stack trace is an essential source of information to locate the bug. The following two predominant categories are Exception Wrapping (10.0%) and Exception Conversion (5.0%), which also possess a correlation between the fix and the stack traces. In only one bug, COMPRESS-31, the exception-related code was not touched for the fix (Keep Throwing category). After a more detailed investigation, we noticed that, despite this bug being in the Defect4j database, its resolution field was marked as “Not A Problem” on Jira, which explains why nothing was done about the exception.
Figure 7: Bugfix intention type distribution in the studied bugs.
In 78.3% of the bugs, the buggy methods are directly reachable from the stack trace, with an average distance of 0.34 method calls. Upon examining the proximity between stack traces and buggy methods, our analysis reveals that 66.67% of the bugs feature at least one buggy method directly present in the stack trace, indicating a zero distance. Furthermore, a total of 78.3% of the bugs contain buggy methods that are reachable from the stack trace (i.e., distance is zero or more), averaging a very short distance from the stack trace methods of 0.34 method calls. In such scenarios, leveraging stack traces to reconstruct the execution path at the time of the exception can be an effective method for pinpointing the bug. We conduct a manual analysis of the remaining 21.7% of the bugs in which the bugs methods are unreachable from the stack trace. We find that, in these cases, the stack traces only contain external entries (i.e., entries referring to external libraries) or the provided stack traces were incomplete. In short, our findings highlight the potential of using stack traces for fault localization.
In 98.3% of the studied bugs, the bugfix intention is directly correlated with the exception in the stack trace (Exception Prevention, Exception Conversion or Exception Wrapping). In addition, 78.3% of the buggy methods are reachable from the stack traces, having an average distance of 0.34 method calls. This shows that the stack traces are a valuable source of information about the bug.
RQ3 - Can we utilize the stack traces to help detect the buggy locations?
Motivation. In RQ1, we found that the existing failing tests rarely matched the actual fault-triggering tests. Without fault-triggering tests, the performance of SBFL techniques is greatly affected. In addition, in RQ2, we found that the stack traces are a valuable source of information about the bug. The intention behind the bugfix is usually related to the exception, and the buggy methods are often a short distance away from the stack traces. Prior studies (Chen et al., 2021a, b) show that developers usually rely on stack traces when investigating for bugs, as they provide essential information about the buggy location. Stack traces, similar to fault-triggering tests, carry contextual and execution information associated with the root causes of bugs. In a way, stack traces can be used to represent the coverage of a failing fault-triggering test. Therefore, in this RQ, we aim to investigate how stack traces can be leveraged to complement test cases in locating bugs.
Approach. To understand how stack trace information can contribute to locating buggy methods, we proposed to evaluate an approach called SBEST (Spectrum-Based localization Enhanced by Stack Traces). SBEST is a fault localization approach based on SBFL that incorporates stack trace information with the test coverage data. This technique applies SBFL principles, but instead of using the failing fault-triggering test coverage in the Ochiai formula, it considers the methods that appear on the stack trace entries as the source of the failure. We define a method that is causing a test failure if it appears on any of the frames in a stack trace. Prior studies (Wong et al., 2014a; Chen et al., 2021b) have also shown that the position of methods in the stack trace can be useful in fault localization. Therefore, we incorporate a Stack Trace (ST) score into our approach. We compare our approach with two baselines. We use Ochiai as the first one because it is shown to be one of the best SBFL techniques and performs very well on real faults (Pearson et al., 2017). In addition, we leverage the position of the stack trace as the second baseline. For this, we rank the top entries in the stack trace as more suspicious. Below, we describe our approach in detail.
Our overall approach, SBEST, consists of the sum of two scores: the SB score and the ST score.
Spectrum Based (SB) Score. The first one, the SB score, is based on an approach designed to use the information from the stack traces in the fault localization process. It is based on the Ochiai formula, and it utilizes existing test coverage information. However, instead of using the failing tests as the hint of the bug location (since there may not be any test failure or the failure is not related to the reported bug, as found in RQ1), we use the tests covering the methods in the stack trace as the fault-triggering tests. We base this change on the fact that the stack trace represents the system execution at the moment of the failure in the same way that the fault-triggering tests do. More specifically, we utilize the methods present on the first five internal entries from the stack traces to select the failing tests that will be applied to Ochiai’s formula. We base this design on the fact that the higher the stack trace entry, the more probable it is to be related to the bug (Wong et al., 2014b).
The selection of failing tests is determined by counting the number of lines each existing test covers in the Stack Trace methods. We first compute the number of covered lines from these methods for each test by applying the following formula:
$ST\_Covered\_lines\_number_{t}=\sum_{m=1}^{5}CL_{m}$ (2)
in which m represents each of the top 5 methods in the Stack Trace, and $CL_{m}$ is the number of lines in the correspondent method that is covered by the test. Then, we select the set of the X highest results by applying the following formula, in which T represents the set of tests and X is a threshold for the number of selected tests:
$\begin{split}T_{\text{failing}}=\{t\in T\mid t\text{ is one of the }X\text{ tests with the highest }\\ ST\_Covered\_lines\_number_{t}\}\end{split}$ (3)
We noticed that selecting a value for X that was too low made the approach often miss buggy locations because it was too focused on just a few methods. On the other hand, a high threshold also generates inaccurate results since the approach is unable to differentiate between the methods in the stack trace. Upon experimentation, we noticed that setting it to 15 gave the best results. Finally, we apply the Ochiai formula (1) using these 15 tests as failing tests and all the remaining as passing tests.
Stack Trace (ST) Score. The SB score described above merges the information in the stack traces with the test coverage to compute the suspiciousness score for a given method. In addition to that, we make use of the ST score used in a previous study (Chen et al., 2021b) in order to boost the stack trace’s impact into the calculated score. To do this, after having computed the SB_score following the approach described before, we sum to it a ST_score calculated in the following manner:
$ST\_score=\begin{cases}\frac{1}{ST\_rank}&\text{if }ST\_rank\leq 10\\ 0.1&\text{if }ST\_rank>10\\ 0&\text{if }\text{method not found}\end{cases}$ (4)
In which ST_rank is the position in which the given method appears in the stack trace after the external entries are removed. For example, if the method appears in the second position, the ST_score would be 1/2=0.5.
Suspiciouness Score. After having both the STCB Ochiai score and the ST Score, we compute the suspiciousness score of a giving method by applying the formula:
$Suspiciouness\_score=STCB\_Ochiai\_score+ST\_score$ (5)
Suspiciouness Rank. After having the final suspiciousness score, we generate the final suspiciousness rankings. The higher the suspiciousness score, the higher the ranking. For instance, if the highest calculated suspiciousness score is 1.0 for the method M8, M8 is going to be the first method in the suspiciousness rank, as it has the highest probability of containing the bug.
Table 5: Fault localization results for all the studied projects. #Bugs represents the total number of studied bugs in each project.
      System (#Bugs)       Technique       Top-1       Top-3       Top-5       MAP       MRR
      Cli (2)       Ochiai       0       0       0       0.0045       0.0045
      Stack Trace       1       1       2       0.6       0.6
      SB_score       0       0       1       0.1333       0.1333
      SBEST       1       1       2       0.6       0.6
      Closure (8)       Ochiai       0       0       0       0.0007       0.0012
      Stack Trace       1       2       2       0.1710       0.2170
      SB_score       0       0       0       0.0003       0.005
      SBEST       1       2       2       0.1462       0.1671
      Codec (1)       Ochiai       0       0       0       -       -
      Stack Trace       1       1       1       0.3333       1
      SB_score       1       1       1       0.3893       1
      SBEST       1       1       1       0.5853       1
      Compress (11)       Ochiai       0       0       1       0.0079       0.023
      Stack Trace       3       6       8       0.4122       0.4425
      SB_score       0       3       3       0.1404       0.1587
      SBEST       3       9       9       0.4620       0.4904
      Csv (1)       Ochiai       0       0       0       -       -
      Stack Trace       0       1       1       0.3333       0.3333
      SB_score       1       1       1       1       1
      SBEST       1       1       1       1       1
      Gson (3)       Ochiai       0       0       0       -       -
      Stack Trace       1       2       2       0.5196       0.5196
      SB_score       1       1       1       0.3338       0.3338
      SBEST       1       2       2       0.5101       0.5101
      JacksonCore (4)       Ochiai       0       1       1       0.0693       0.0836
      Stack Trace       1       1       1       0.25       0.25
      SB_score       0       0       0       0.0409       0.0457
      SBEST       0       1       1       0.1609       0.1634
      JacksonDatabind (10)       Ochiai       0       0       0       0.0005       0.0005
      Stack Trace       1       2       3       0.1686       0.2446
      SB_score       0       1       1       0.0125       0.0212
      SBEST       1       3       3       0.0359       0.0777
      Jsoup (8)       Ochiai       0       0       0       0       0
      Stack Trace       3       3       5       0.4410       0.4444
      SB_score       0       0       0       0.0214       0.0318
      SBEST       3       4       4       0.4070       0.4279
      JxPath (1)       Ochiai       0       0       0       -       -
      Stack Trace       1       1       1       0.3333       0.3333
      SB_score       0       0       0       0.0625       0.0625
      SBEST       1       1       1       1       1
      Lang (4)       Ochiai       0       0       0       0.0008       0.0008
      Stack Trace       1       3       3       0.4375       0.4375
      SB_score       0       1       1       0.1411       0.1411
      SBEST       2       3       3       0.5835       0.5835
      Math (3)       Ochiai       0       0       0       0.0002       0.0001
      Stack Trace       2       2       2       0.2215       0.4583
      SB_score       0       0       0       0.0428       0.0292
      SBEST       2       2       2       0.1372       0.3472
      Mockito (2)       Ochiai       0       0       0       0.0060       0.0281
      Stack Trace       0       1       1       0.0238       0.1667
      SB_score       0       1       1       0.0371       0.25
      SBEST       0       1       1       0.0371       0.25
      Time (2)       Ochiai       0       0       0       0.0003       0.0003
      Stack Trace       0       1       2       0.2917       0.2917
      SB_score       1       1       1       0.5001       0.5001
      SBEST       0       1       1       0.3333       0.3333
      Total (60)       Ochiai       0       1       2       0.00902       0.01421
      Stack Trace       16       27       34       0.32406       0.40992
      SB_score       4       10       11       0.20396       0.26517
      SBEST       17       32       33       0.42846       0.49647
Results. The Stack Traces ranking alone locates 34 out of the 60 bugs in the Top-5. Table 5 presents the metrics results for the 2 baselines as well as for the SB_score alone and the SBEST approach for all the systems. We can see that the worst-performing technique is Ochiai, which, as discussed before, locates only 2 bugs in Top-5. The best-performing approach is SBEST, but the Stack Traces ranking is a close second. The Stack Trace ranking locates 16 bugs in Top-1, 27 in Top-3 and 34 in Top-5. Surprisingly, the Stack Traces alone provide very good results, locating more than half of the bugs (56.67%) within the Top 5. This shows that the stack traces are a very important source of information about the bug location and should be prioritized when it comes to studying FL in crash reports. The SB_score is able to locate some bugs that the Stack Trace ranking does not locate, but in general, it does not perform very well, identifying only 18.33% of the buggy methods in the Top 5. We can see, however, that the SB_score is substantially improved by the addition of the ST_score. The resultant technique SBEST performs slightly better than the Stack Traces in the Top 1 and Top 3, locating 17 and 32 bugs, respectively, and locates only one bug less in the Top 5, totalling 33 bugs spotted within the Top 5 rank. This technique and the Stack Traces ranking take turns as the best localization approach for each project. For example, the Stack Trace ranking is better overall for the JacksonCore project, while SBEST has the top performance on the Lang bugs. In total, SBEST locates 17 bugs in Top-1, 32 in Top-3 and 33 in Top-5. The average MAP and MRR for SBEST are 0.42846 and 0.49647, respectively, representing a 32.22% improvement on the MAP and a 17.43% in the MRR when compared to the Stack Traces. These results show that the stack traces can complement existing passing test information to help locate the bug root causes.
Overall, out of 60 studied bugs, SBEST successfully located 17 bugs in Top-1, 32 in Top-3 and 33 in Top-5. In addition, it achieves an improvement of 32.22% on MAP and 17.43% on MRR when compared to the Stack Trace ranking.
6 Discussion and Implication of Our Findings
On the use of stack traces in the absence of failing tests. From the results obtained in Section 5, we observe that the majority of the crash report bugs do not contain fault-triggering tests. In such cases, the test failure information is not available for the fault localization process, causing techniques such as SBFL to fail. The stack traces, on the other hand, demonstrate to be deeply correlated to the bug cause. We found that, in 66% of the bugs, at least one buggy method is directly listed in the stack trace. Even when this is not the case, the buggy methods are usually just a few calls away from it. In addition, the most common bugfix intention is to prevent the exception in the stack trace from happening again, which highlights the big association between them.
Leveraging stack trace rankings for enhanced fault localization. In the real world, test failures are not always present. Especially when it comes to production-phase bugs, it is important to look for alternative sources of information for fault localization. Making use of the data available in logs and stack traces shows to be a good path to take in such scenarios. Even though previous studies have used stack traces for Fault Localization, many are based on information retrieval (IRFL) approaches (Zhou et al., 2012; Saha et al., 2013; Lam et al., 2017). IRFL techniques treat all the content in the bug report, including the stack traces, as textual information, therefore missing important context such as the ranking of the stack traces. Our results show that the stack traces ranking alone was able to locate more than half of the bugs within Top-5 and, thus, should be better utilized.
Encouraging the integration of stack traces to complement execution information. SBEST, our approach combining the stack trace ranking with the code coverage, shows promising results. The improvement in the metrics implies that the coverage information was able to enhance the stack trace ranking, aiding in obtaining a better overview of the system execution in the moment of failure. Our main goal when studying this technique was to gain a deeper understanding of the scenario of these bugs and how each piece of information available can help with fault localization. We believe that this is just the first step and that future research can benefit from the aforementioned findings to build more intricate techniques. In addition, we make all the data from this study available at Zenodo [^1] https://zenodo.org/records/11062413, including the detailed code coverage and test results of the bug report commits that took us 100 hours to obtain. We think that this data will be useful for future research on subjects such as fault localization, automatic bug fixing, test generation, etc. Our data does not contain post-bug-fixing development knowledge, which will allow researchers to tailor tools more aligned with the real-world scenario.
7 Threats to Validity
In this section, we discuss the threats to validity related to this study.
7.1 External Validity.
Threats to external validity relate to how generalizable our findings are. To minimize this threat, we conduct our studies in 15 systems from Defects4j, a famous framework of real bugs used in multiple other studies. These systems are widely used and vary in size and number of tests. Even though every system is written in Java, our method is not restricted to Java-based systems. In addition, despite the low percentage of bugs with stack traces in Defects4j, this aligns with the percentage reported in other systems (Chen et al., 2021a). Moreover, we opted not to use the buggy version made available in defects4j but the code available at the moment of the bug report creation in order to better emulate the scenario in which the developers investigate the bug. Finally, even though we opted for applying only Ochiai when analyzing SBFL’s efficiency when applied to crash report bugs, it is shown to be one of the best SBFL techniques and performs very well on real faults (Pearson et al., 2017).
7.2 Internal validity.
Threats to internal validity relate to the extent to which the results of the study can be attributed solely to the experimental treatments and not to flaws in the experimental design. In this study, we only analyze Defects4j bugs. Although Defects4j is widely used (Lutellier et al., 2020; Chen et al., 2021b; Li et al., 2021; Ye et al., 2022), the findings might be different in other systems, specially non-Java systems. Another threat to internal validity is the way in which we selected the bug reports with stack traces. We collect stack traces in the description and comments, which is the typical place in which they are added. However, in some cases, developers may also add stack traces via attachments. Nonetheless, this is very rare, representing less than 1% of the cases based on a previous study from Chen et al. (2021a).
7.3 Construct validity.
Construct validity refers to how well the study’s procedures and metrics accurately capture the concepts they intend to investigate. We use three evaluation metrics in our study: Top K, MAP, and MRR. These metrics are commonly used in FL and have been used in many previous studies (Wen et al., 2021; Chen et al., 2021b; Xia et al., 2016; Kochhar et al., 2016).
8 Conclusion
Spectrum-based fault Localization (SBFL) approaches are highly recognized due to their accuracy, efficiency, and simplicity. However, their effectiveness depends on the availability of fault-triggering failing tests. In the lack of such test failures, alternative approaches should be considered. In this paper, we research using stack traces for SBFL in the absence of failing tests. We study 60 crash reports from the Defects4J benchmark. We find that most of the studied crash reports do not contain fault-triggering failing tests. This results in very low efficiency when using traditional SBFL. On the other hand, when it comes to the stack traces, we find that in 98.3% of the studied bugs, the bugfix intention is directly correlated with the exception in the stack trace. In addition, in 78.3% of the bugs, the buggy method is directly reachable from the stack trace, with an average distance of only 0.34 method calls. This shows that the buggy methods are usually very close to the stack traces when it comes to the execution call graph. On top of that, our results show that even without any advanced technique, the stack traces alone provide a good indication of the buggy locations, being capable of locating more than half of the bugs on Top-5. Finally, we develop a simplified SBFL method called SBEST that uses the stack trace information in place of the failing tests and thus integrates it into the coverage information to perform the SBFL. SBEST was able to locate 17 bugs in Top-1, 32 in Top-3 and 33 in Top-5, which represents an improvement of 32.22% on MAP and 17.43% on MRR when compared to the Stack Trace ranking. The enhancement of the metrics suggests that the coverage information improved the stack trace ranking and helped the technique gain a clearer picture of the system’s execution at the time of failure. We believe that future studies can benefit from the findings of this research to develop more intricate techniques. In addition, all the data that we have made available can help studies in areas such as automatic bug fixing, test generation, and, of course, fault localization. As potential future work, we plan to expand our analysis by increasing our bugs dataset utilizing other sources beyond Defects4j.

### v2

@@
-An Ran Chen
+
 Abstract
-Bug fixing is a crucial task in software maintenance to hold user trust.
-Although various automated fault localization techniques exist, they often require specific conditions to be effective.
-For example, Spectrum-Based Fault Localization (SBFL) techniques need at least one failing test to identify bugs, which may not always be available.
-Bug reports, particularly those with stack traces, provide detailed information on system execution failures and are invaluable for developers.
-This study focuses on utilizing stack traces from crash reports as fault-triggering tests for SBFL.
-Our findings indicate that only 3.33% of bugs have fault-triggering tests, limiting traditional SBFL efficiency.
-However, 98.3% of bugfix intentions align directly with exceptions in stack traces, and 78.3% of buggy methods are reachable within an average of 0.34 method calls, proving stack traces as a reliable source for locating bugs.
-We introduce a new approach, SBEST, that integrates stack trace data with test coverage to enhance fault localization.
-Our approach shows a significant improvement, increasing Mean Average Precision (MAP) by 32.22% and Mean Reciprocal Rank (MRR) by 17.43% over traditional stack trace ranking methods.
-Keywords:
-bug report, stack trace, empirical study, fault localization, software maintenance.
+Fault localization is a critical step in software maintenance.
+Yet, many existing techniques, such as Spectrum-Based Fault Localization (SBFL), rely heavily on the availability of fault-triggering tests to be effective.
+In practice, especially for crash-related bugs, such tests are frequently unavailable.
+Meanwhile, bug reports containing stack traces often serve as the only available evidence of runtime failures and provide valuable context for debugging.
+This study investigates the feasibility of using stack traces from crash reports as proxies for fault-triggering tests in SBFL.
+Our empirical analysis of 60 crash-report bugs in Defects4J reveals that only 3.33% of these bugs have fault-triggering tests available at the time of the bug report creation.
+However, 98.3% of bug fixes directly address the exception observed in the stack trace, and 78.3% of buggy methods are reachable within an average of 0.34 method calls from the stack trace.
+These findings underscore the diagnostic value of stack traces in the absence of failing tests.
+Motivated by these findings, we propose SBEST, a novel approach that integrates stack trace information with test coverage data to perform fault localization when fault-triggering tests are missing.
+SBEST shows an improvement, with a 32.22% increase in Mean Average Precision (MAP) and a 17.43% increase in Mean Reciprocal Rank (MRR) compared to baseline approaches under the scenario where fault-triggering tests are absent.
 1 Introduction
-Locating and fixing software bugs is an essential yet frequently expensive and time-consuming task.
-Bug-fixing activities not only delay the project’s progress but also increase the software’s overall cost, making them a critical part of software maintenance.
-Considering the complexity and rapid evolution of modern software systems, the ability to efficiently and accurately locate bugs is more important than ever.
-To address this challenge, previous research has proposed numerous different automated fault localization techniques (Abreu et al., 2007; Jones and Harrold, 2005; Naish et al., 2011; Abreu et al., 2009; Wong et al., 2013; Papadakis and Le Traon, 2015; Hong et al., 2017; Papadakis and Le Traon, 2012; Wang et al., 2022; Cui et al., 2020), among which Spectrum-Based Fault Localization (SBFL) approaches are prominent due to its high accuracy (de Souza et al., 2016).
-SBFL techniques leverage information derived from test executions — specifically, the spectrum of code elements executed during passing and failing tests — to calculate the probability of every code element being faulty (i.e.
-suspiciousness score).
-The effectiveness of SBFL highly depends on the availability of failing test cases that uncover the bug (i.e., fault-triggering tests).
-However, studies indicate that in real-world scenarios, especially in continuous deployment and integration environments, failing tests may not always be present or identified when bugs are reported (Chen et al., 2023; Kabadi et al., 2023).
-This absence of fault-triggering tests significantly impacts the applicability of SBFL, as it depends on it to distinguish between faulty and non-faulty code.
-In the absence of failing tests, alternative approaches must be considered.
-Exception stack traces, commonly included in crash reports, provide dynamic execution information that can be extremely valuable for debugging.
-Studies show that developers usually rely on them when investigating bugs (Schroter et al., 2010; Zimmermann et al., 2010).
-In addition, previous research (Chen et al., 2021a) shows that stack traces provide valuable information on the cause of the bugs and their location in the source code.
-These stack traces, which outline the sequence of method calls leading up to an exception, offer implicit clues about where a bug may reside, providing a fallback localization method in environments where traditional test-based approaches falter.
+Locating and fixing software bugs is an essential but often expensive and time-consuming task.
+Bug-fixing activities delay project progress and increase overall maintenance costs.
+As modern software systems grow in size and complexity, efficiently and accurately localizing faults has become increasingly critical.
+To address this challenge, previous research has proposed numerous different automated fault localization techniques (Abreu et al., 2007; Jones and Harrold, 2005; Naish et al., 2011; Abreu et al., 2009; Wong et al., 2013; Papadakis and Le Traon, 2015; Hong et al., 2017; Papadakis and Le Traon, 2012; Wang et al., 2022; Cui et al., 2020).
+Among these, Spectrum-Based Fault Localization (SBFL) stands out due to its conceptual simplicity, low computational cost, and exemplary performance in many scenarios (de Souza et al., 2016).
+SBFL techniques leverage information derived from test executions, specifically the spectrum of code elements executed during passing and failing tests, to calculate the probability of every code element being faulty (i.e., suspiciousness score).
+Although SBFL is no longer considered state-of-the-art in terms of absolute accuracy, it remains highly relevant and widely adopted today.
+For example, prior learning-based fault localization techniques (Li et al., 2019; Sohn and Yoo, 2017) use SBFL scores as core input features for model training.
+More recently, LLM-based approaches like AutoCodeRover (Zhang et al., 2024) and LLAMAO (Yang et al., 2024) also rely on SBFL outputs to guide code analysis and method ranking.
+Beyond fault localization, SBFL plays a central role in various downstream software engineering tasks, including test case prioritization (Li et al., 2007; Srivastava, 2008), and automated program repair (APR) (Motwani and Brun, 2023; Kuma et al., 2020; Lou et al., 2020).
+The effectiveness of SBFL highly depends on the availability of failing test cases that uncover the bug, known as fault-triggering tests (Chen et al., 2022).
+A fault-triggering test is a test case that fails specifically because it executes the buggy code responsible for the fault, and passes once the bug is fixed.
+However, studies indicate that in real-world scenarios, especially in continuous deployment and integration environments, such tests may not always be present or identified when bugs are reported (Rafi et al., 2025; Kabadi et al., 2023).
+The absence of fault-triggering tests can significantly impact the applicability of SBFL, as it relies on them to distinguish between faulty and non-faulty code.
+In such situations, developers often rely on Crash Reports, a specific type of bug report that documents unexpected program terminations along with diagnostic details such as exception messages and stack traces.
+A crash report typically includes a title, description, and the stack trace that captures the program’s execution path leading to the failure (Chen et al., 2021a).
+For example, Figure 1 shows a crash report from the Defects4J 2.0 dataset (COMPRESS-181), which describes an IOException triggered when reading TAR files with symbolic links.
+The report includes the relevant stack trace but no fault-triggering tests at the time, making SBFL unable to prioritize the faulty locations.
+This example highlights how stack trace sequences of method calls preceding an exception can serve as an alternative source of execution information when traditional test-based approaches are not applicable (Schroter et al., 2010; Wu et al., 2014).
+Figure 1: Example bug report (COMPRESS-181).
+The bug report addresses a problem in reading TAR files that contain symlinks, which cause an IOException to be thrown.
+The report contains the stack trace of the error (gray box), and an example of TAR file (simple-aix-native-tar.tar) that triggers the bug is attached.
+When this bug report was created, there were no failing tests.
 In this paper, we address the challenges of fault localization in scenarios where traditional fault-triggering test data is scarce.
-We conduct our study on the Defects4J 2.0 dataset (Just et al., 2014b), a widely recognized resource for research in fault localization and program repair.
-However, prior studies (Kabadi et al., 2023; Chen et al., 2023) have reported that some tests in Defects4J were added post hoc, following the reporting of bugs, which could bias the bug localization process.
-To avoid this issue and better simulate the state of the codebase at the time developers first address a bug, we utilize the commit data immediately preceding the bug report rather than relying directly on the buggy commits provided by Defects4J.
-Our initial analysis of the bugs reveals a critical limitation: only 3.33% of crash reports, which are bug reports that contain stack traces, possess fault-triggering tests.
-Even for bug reports without stack traces, only 10.19% of the bugs have fault-triggering tests.
+We conduct our study on the Defects4J 2.0 dataset (Just et al., 2014b), a widely recognized benchmark for research in fault localization and program repair.
+Prior studies (Kabadi et al., 2023; Rafi et al., 2025) report that some Defects4J tests were added post hoc, potentially biasing localization.
+To avoid this, we analyze the commit preceding the bug report to better reflect the state of the code repository that developers saw at the time.
+Our initial analysis of the bugs reveals that only 3.33% of crash reports, which are bug reports that contain stack traces, possess fault-triggering tests.
+Even for bug reports without stack traces, only around 10% of the bugs have fault-triggering tests.
 This low incidence of failing tests challenges the efficacy of Spectrum-Based Fault Localization (SBFL) methods, which rely heavily on test outcomes to identify bug locations.
-To overcome this limitation, we propose leveraging the stack traces—rich in contextual and execution information about the bug’s root causes—as a substitute for the failing tests.
+To overcome this limitation, we propose leveraging stack traces, which are rich in contextual and execution information about the bug’s root causes, as a substitute for the failing tests.
 We introduce a novel approach, Spectrum-Based Localization Enhanced by Stack Traces (SBEST), which integrates stack trace analysis within the SBFL framework.
-Our method significantly improves the accuracy of fault localization by utilizing the comprehensive code execution information allied with the exception stack traces.
+Our method improves fault localization accuracy in scenarios where fault-triggering tests are unavailable by combining coverage information with the execution context provided by exception stack traces.
 This integration enables a more detailed understanding of system behaviour at the moment of failure, thereby addressing the gaps left by the unavailability of failing tests.
@@
   • RQ1: Are the test failures related to the bug in crash reports?
-We study the presence of fault-triggering tests in crash reports (i.e., bug reports that contain stack traces) and their impact on the SBFL results.
-We find that only 3.33% of the crash report bugs contain fault-triggering tests, which profoundly degrades the SBFL efficiency in locating these bugs.
-The limited execution data underscores the importance of incorporating stack traces into the SBFL process, enhancing the potential to accurately identify bug locations in the absence of traditional test failures.
+We examine whether the test failures observed at the bug report commit are truly fault-triggering, that is, caused by the bug described in the crash report.
+Our analysis shows that only 3.33% of crash-report bugs have such fault-triggering tests, limiting the applicability of SBFL.
+This finding underscores the need to augment SBFL with alternative sources of execution information, such as stack traces, especially when traditional failing tests are unavailable.
   • RQ2: What is the relationship between stack traces and buggy location?
@@
 In addition, 78.3% of the buggy methods are directly reachable from the stack traces, with an average distance of 0.34 method calls.
-This indicates that leveraging the stack traces to reconstruct the execution path at the time of the exception can be an effective method for pinpointing the bugs.
+This indicates that leveraging the stack traces to reconstruct the execution path at the time of the exception can be an effective method for pinpointing bugs.
   • RQ3: Can we utilize the stack traces to help detect the buggy locations?
-Given the previous findings, we propose a new approach, SBEST, that integrates the stack trace information into Spectrum-Based Fault Localization.
-We compare our approach with two baselines: the Ochiai results and the position of the methods in the stack trace.
-Ochiai proves to be the least effective technique by a significant margin.
-On the other hand, we find that the stack traces alone can locate 34 out of the 60 bugs in the Top-5.
-This reinforces that the stack traces are a vital source of information about the bug location and should be prioritized in FL.
-In addition, SBEST achieves an improvement of 32.22% on MAP and 17.43% on MRR compared to the Stack Trace ranking.
-The enhancement of metrics suggests that the incorporation of coverage information effectively improved the ranking of stack traces, thereby providing the tool with a more comprehensive view of system execution at the time of failure.
-In summary, this paper presents a comprehensive study that evaluates the effectiveness of combining stack traces with test coverage data.
-Our findings indicate that by incorporating stack traces into the SBFL model, we can significantly improve bug localization, thereby addressing one of the key challenges in automated debugging processes.
-We make the data publically available online for replication and encourage future research in this area (sbe, 2024).
+We propose SBEST, a technique that leverages stack traces as a proxy for fault-triggering tests within the SBFL framework.
+To evaluate SBEST, we compare it against four baselines: Ochiai, Stack Trace ranking, SB_score (stack trace + Ochiai), and Call Frequency-Based Fault Localization.
+Among the baselines, Stack Trace alone ranks 34 out of 60 bugs in the Top-5, while Ochiai identifies only 2.
+SBEST localizes 33 bugs in the Top-5, with MAP and MRR values 32.2% and 17.4% higher than the Stack Trace baseline, respectively.
+In summary, this paper presents a comprehensive study evaluating the effectiveness of combining stack traces with test coverage data.
+Our findings indicate that incorporating stack traces into the SBFL model can enhance bug localization in scenarios where fault-triggering tests are unavailable, addressing an important challenge in automated debugging.
+We make the data publicly available online for replication and encourage future research in this area (Rafi et al., 2024).
 Paper Organization.
-Section 2 details the background and the motivation behind this study.
 Section 2 surveys the related work.
@@
 Section 7 concludes the paper and discusses future work.
-2 Background and Motivation
-Bug Reports contain important information for developers to fix the issues that users or other developers encounter in the software (Chen et al., 2021a).
-A bug report typically includes a title, a description, attached files, and comments.
-For example, Figure 1 shows a bug report from the Defects4j 2.0 dataset.
-The title offers a concise one-line description of the problem.
-The description provides developers with the necessary details to investigate the bug, such as the stack traces illustrating the system execution details.
-When additional information is required, users and developers use the comments and attached files sections to further communicate on the bug.
-In Figure 1 bug COMPRESS-181, a problem occurs when the system reads TAR files with symlinks.
-In particular, the description provides developers with a stack trace that illustrates an IOException is being thrown.
-The description also provides the affected nightly builds, which show no test failures.
-Figure 1: An example of a bug report (COMPRESS-181).
-The bug report addresses a problem in reading TAR files that contain symlinks, which cause an IOException to be thrown.
-The report contains the stack trace of the error (gray box), and an example of TAR file (simple-aix-native-tar.tar) that triggers the bug is attached.
-When this bug report was created, there were no failing tests.
-To assist developers in debugging such bugs, prior studies have proposed Spectrum-Based Fault Localization (SBFL) techniques (Chen et al., 2022; Lou et al., 2021; Baudry et al., 2006).
-SBFL relies on the information from failing tests to identify potentially faulty code.
-Hence, SBFL assumes that test failures that are relevant to the bug exist.
-SBFL works by comparing the code execution of passing and failing tests to identify their differences.
-The potential problematic code is then identified based on these differences.
-SBFL is widely studied in prior studies for its simplicity and efficiency (Santelices et al., 2009; Pearson et al., 2016).
-However, when there is no failing test, SBFL can not distinguish buggy code elements.
-For instance, in the case of bug COMPRESS-181 (Figure 1), in which there are no failing tests, SBFL can not help prioritize the faulty locations.
-This is because, in the absence of failing tests, there is no coverage spectrum between passing and failing executions.
-Thus, it becomes impossible to prioritize any code element for investigation, as SBFL relies on failing executions as a guide to identify where faults might be present.
-In this case, all the code locations receive the exact same suspiciousness score of zero, and the bug cannot be located.
-In this paper, we investigate in detail the characteristics of Defects4j 2.0 crash reports (i.e., bug reports that contain stack traces).
-We dissect information such as the test failures, coverage, and stack traces by conducting a quantitative and qualitative analysis of the data and propose a new approach to address the limitations found in SBFL in such scenarios.
 2 Related Work
 2.1 Fault Localization
-Automated fault localization methods are a recurrent theme in the literature, being Spectrum-Based Fault Localization (SBFL) and Mutation Based Fault Localization (MBFL) two of the most famous representatives.
+Automated fault localization methods are a recurrent theme in the literature, being Spectrum-Based Fault Localization (SBFL) and Mutation-Based Fault Localization (MBFL) being two of the most famous representatives.
 SBFL is a very consolidated class of Fault Localization techniques.
-These approaches use the test information to localize the bugs, and they require at least one failing test to work.
+These approaches utilize test information to localize bugs, and they require at least one failing test to be effective.
 Ochiai (Abreu et al., 2007) is one of the most famous SBFL techniques.
-Other representatives of this class are Tarantula (Jones and Harrold, 2005), Op2 (Naish et al., 2011), BARINEL (Abreu et al., 2009) and DStar (Wong et al., 2013): similar approaches based on different formulas to compute the suspiciousness score.
-Given that SBFL techniques are light weighted and fast, they are also applied to other kinds of techniques such as interactive fault-localization (Gong et al., 2012) and program repair (Ye et al., 2022; Le Goues et al., 2012).
+Other representatives of this class are Tarantula (Jones and Harrold, 2005), Op2 (Naish et al., 2011), BARINEL (Abreu et al., 2009), and DStar (Wong et al., 2013): similar approaches based on different formulas to compute the suspiciousness score.
+Given that SBFL techniques are lightweight and fast, they are also applied to other types of methods, such as interactive fault localization (Gong et al., 2012) and program repair (Ye et al., 2022; Le Goues et al., 2012).
+Information Retrieval-based Fault Localization (IRFL) approaches aim to localize buggy code by treating bug reports as textual queries and comparing them against the source code using similarity metrics.
+Early techniques, such as BugLocator (Zhou et al., 2012), enhanced traditional IR techniques by combining TF-IDF-based scoring with historical bug-fix data to improve the ranking of relevant files.
+Subsequent work, such as BLUiR (Saha et al., 2013), further advanced this by leveraging structured information from source code, including method and class names, to improve retrieval accuracy over treating code as flat text.
+In addition, BRTracer (Wong et al., 2014) improved IR-based fault localization by segmenting source files and boosting files referenced in stack traces.
+Additionally, hybrid approaches, such as the one proposed by Lam et al.
+(2017), combine deep learning (e.g., CNNs) with IR signals to better capture the semantic relationships between bug reports and source code, outperforming purely IR-based baselines.
+Supervised IR-based models such as STMLocator+ (Wang et al., 2020) further enhanced accuracy by integrating textual and semantic similarity, metadata, and stack trace information.
 MBFL techniques (Papadakis and Le Traon, 2015; Moon et al., 2014; Hong et al., 2017; Papadakis and Le Traon, 2012; Wang et al., 2022) are based on mutation analysis.
@@
 Despite presenting promising results, this kind of technique is very costly and time-consuming since it requires multiple executions of the test suite.
-Strategies for reducing this overhead and improving the general accuracy have been studied, such as ways prioritizing the mutants (Liu et al., 2018), prioritizing the tests executed for each mutant (de Oliveira et al., 2018) and a mix of spectrum and mutant-based approaches (Cui et al., 2020).
+Strategies for reducing this overhead and improving the general accuracy have been studied, such as ways of prioritizing the mutants (Liu et al., 2018), prioritizing the tests executed for each mutant (de Oliveira et al., 2018), and a mix of spectrum and mutant-based approaches (Cui et al., 2020).
 One common point between SBFL and MBFL approaches is that both types of techniques require failing tests to work.
-However, this might be not always the case in real life.
+However, this might not always be the case in real life.
+Call Frequency-Based Fault Localization (CFBFL) (Vancsics et al., 2021) is another dynamic approach that does not rely on coverage spectra but instead ranks methods based on how frequently they are executed across tests.
+Unlike simple count-based methods, CFBFL considers the number of times a method appears in unique call stack contexts during failing executions, filtering out noise from loops and repetitive calls.
+A related approach by  Beszedes et al.
+(2020) improves SBFL by analyzing function call chains, computing suspiciousness over call paths, and then attributing it to individual functions.
+Both techniques enhance fault localization by leveraging richer execution context beyond binary coverage.
 Most of the mentioned studies use Defects4J to evaluate the results.
-Defects4j is one of the most famous benchmarks for test-based fault localization research.
-Nonetheless, recent studies (Kabadi et al., 2023; Chen et al., 2023) show that often in Defects4j, tests that were created after the bug-report creation are added to the buggy patch and thus contain developer knowledge about the bug investigation process.
-These tests were artificially appended to the buggy commit in Defects4j to simulate the scenario of a bug with test failures.
-This fact indicates that many fault localization techniques that were evaluated in Defects4j might have very different results when applied to real-world scenarios.
+Defects4J is one of the most widely used benchmarks for fault localization techniques that rely on test coverage to identify suspicious code elements.
+Nonetheless, recent studies (Kabadi et al., 2023; Rafi et al., 2025) show that often in Defects4J, tests created after the bug report are added to the buggy patch and thus contain developer knowledge about the bug investigation process.
+These tests were artificially appended to the buggy commit in Defects4J to simulate a scenario with a bug and test failures.
+This fact suggests that many fault localization techniques evaluated in Defects4J may yield significantly different results when applied to real-world scenarios.
 2.2 Bug Reports and Stack Traces
-The bug report quality is a common topic in previous research.
-Knowing which parts of the report are more important in the bug-fixing process is very important to guide new studies on fault localization.
-Our focus on the stack trace information is supported by some previous results.
-In the work by Bettenburg et al.
-(2008), the authors conduct a survey to understand what makes a good bug report from the point of view of the developers.
-They found that information about the steps to reproduce the bug, stack traces, and test cases are considered very helpful.
-In the work by Schroter et al.
-(2010), the authors analyze the intersection between the methods in the stack trace and the methods modified in the bugfix commit.
-They found out that more than 47% of the stack traces extracted from bug reports contained at least one buggy method.
-An empirical study from Chen et al.
-(2021a) analyzes the logs’ (stack traces and log snippets) importance in crash reports.
-By extensively analyzing 1,561 bug reports with logs, they concluded that logs often provide helpful information about where the bug is located.
-They found out that in 73% of the bug reports, there is an overlap between the classes in the logs and the fixed classes.
-4 Data Collection and Case Study Setup
+Bug report quality has long been a central topic in software engineering research.
+Understanding which elements of a bug report are most helpful in identifying and resolving faults can help guide the development of more effective fault localization techniques.
+One of the most consistently valuable components of bug reports is the stack trace, particularly in the context of crashes and runtime errors.
+Bettenburg et al.
+(2008) conducted a developer survey to determine what makes a bug a helpful report.
+Their results highlight the importance of detailed reproduction steps, stack traces, and associated test cases.
+These elements help developers quickly understand and reproduce the issue, making stack traces especially valuable for debugging.
+Schroter et al.
+(2010) further investigated the connection between stack traces and the actual faults by analyzing whether methods in the reported stack traces were modified during the corresponding bug fix.
+Their findings revealed that in over 47% of the cases, at least one buggy method appeared in the stack trace, suggesting that stack traces often point directly to fault-relevant areas of the code.
+Building on this, Chen et al.
+(2021a) performed a large-scale empirical study of 1,561 bug reports containing logs, including stack traces and log messages.
+They found that in 73% of the cases, there was an overlap between the classes mentioned in the logs and the classes involved in the bug fix.
+This reinforces the view that logs and especially stack traces provide strong signals for localizing faults.
+Several techniques have been proposed that leverage stack traces more directly for automated fault localization.
+For example, Jiang et al.
+(2012) use stack traces in conjunction with program slicing to localize null pointer exceptions better.
+Gong et al.
+(2014) generates stack traces to aid in crash fault localization and demonstrates their approach’s effectiveness by identifying 64% of crashing faults in Firefox 3.6.
+CrashLocator (Wu et al., 2014) introduced a scalable method for leveraging stack traces across multiple crash reports.
+It clusters similar crash reports into buckets corresponding to the same fault and cross-analyzes the slightly varying stack traces within each bucket.
+The approach expands stack traces using static call graph traversal and ranks methods based on their frequency within the cluster, rarity across other clusters, and proximity to the reported crash point.
+This design allows CrashLocator to combine statistical signals with structural code context, yielding highly accurate fault predictions even in large-scale systems.
+Together, these studies and tools highlight the value of stack traces not only as debugging aids for developers but also as key artifacts that can be mined and augmented for automated fault localization.
+Comparison with Closely Related Studies.
+While prior studies have demonstrated the usefulness of stack traces for fault localization, their objectives and assumptions differ from ours.
+CrashLocator (Wu et al., 2014) clusters multiple crash reports and expands stack traces using static call graphs to rank methods by their frequency and proximity within crash clusters.
+In contrast, SBEST focuses on individual crash reports and integrates stack-trace-derived execution paths directly into the Spectrum-Based Fault Localization (SBFL) framework, treating stack-trace methods as proxies for fault-triggering tests.
+Call Frequency-Based Fault Localization (CFBFL) (Vancsics et al., 2021) extends Spectrum-Based Fault Localization by weighting coverage information with the frequency of method occurrences in stack traces collected during test executions.
+Information Retrieval (IR)-based techniques such as BugLocator (Zhou et al., 2012) and STMLocator+ (Wang et al., 2020) rely mainly on textual similarity between bug reports and source code, occasionally incorporating stack traces as supplementary input.
+In contrast, SBEST leverages stack traces extracted from crash reports rather than from executed tests and integrates them with coverage data to perform fault localization even when fault-triggering tests are missing.
+3 Data Collection and Study Design
 In this section, we start by describing the data collection process and presenting our dataset. After this, we detail the evaluation metrics.
@@
 However, there are some limitations in Defects4J.
-Recent studies (Kabadi et al., 2023; Chen et al., 2023) found that many bugs in Defects4J v1.0 contain tests from the “future” (i.e., added by developers after the bugs were fixed) in their designated buggy commits.
+Recent studies (Kabadi et al., 2023; Rafi et al., 2025) found that many bugs in Defects4J v1.0 contain tests from the “future” (i.e., added by developers after the bugs were fixed) in their designated buggy commits.
 This is problematic because the test coverage may contain developers’ knowledge of the bug, which can cause noise and bias in the result of the downstream research.
@@
 We refer to this commit as a bug report commit to differentiate it from the buggy commit provided by Defects4J.
-Due to the above-mentioned issues in Defects4J, our next step is to collect the test results, and detailed test coverage for the bug report commits.
+Due to the above-mentioned issues in Defects4J, our next step is to collect the test results and detailed test coverage for the bug report commits.
 To do that, we utilized GZoltar (Campos et al., 2012).
@@
 In this work, we target user-reported bugs with stack traces reported, i.e., crash-reporting bugs.
-Among the 15 projects, there are 89 crash-reporting bugs (i.e.
-bugs with stack traces in their bug reports), which represents 11.08% of the total set.
+Among the 15 projects, there are 89 crash-reporting bugs (i.e., bugs with stack traces in their bug reports), which represent 11.08% of the total set.
 Of these 89 bugs, we were able to compile and collect the test execution information for 63 bugs.
@@
 In total, we study 60 bugs.
-Despite the low percentage of crash reports found in Defects4j, they match with the percentage of bugs with stack traces in other projects (Chen et al., 2021a).
+Despite the low percentage of crash reports found in Defects4J, they match the percentage of bugs with stack traces in other projects (Chen et al., 2021a).
 3.2 Evaluation Metrics
-In this paper, we conduct a series of experiments on locating bugs using stack traces.
-In order to evaluate and compare the results, we utilize a set of well-consolidated metrics in the fault localization field.
-Previous studies suggest that performing fault-localization at the file level lacks precision (Kochhar et al., 2016) while opting for the statement-level granularity can cause the approach to miss important code context (Parnin and Orso, 2011).
-Hence, in this paper, we conduct our analysis at the method level.
+Previous studies suggest that performing fault localization at the file level lacks precision (Kochhar et al., 2016), while opting for statement-level granularity can cause the approach to miss important code context (Parnin and Orso, 2011).
+A survey of developers (Kochhar et al., 2016) further indicates that method-level granularity is the most preferred approach for fault localization tasks.
+Hence, we conduct our analysis at the method-level in this paper by following prior studies (B.
+Le et al., 2016; Kang et al., 2024b; Benton et al., 2020; Lou et al., 2021).
 Mean Average Precision (MAP).
-MAP is a metric that considers the rank of all the buggy methods in a given project.
-It is calculated by taking the mean of the Average Precision (AP) across all these bugs, where the following formula defines the AP, and $m$ is the set of buggy files from a given bug report.
-$AP=\frac{\sum_{i=1}^{m}{i}/{Pos(i)}}{m}$
+MAP is a popular metric for evaluating the effectiveness of ranking systems, such as fault localization (FL) techniques (Sohn and Yoo, 2017; Chen et al., 2021b; Hirsch and Hofer, 2023).
+It is the mean of the average precision of all faults.
+It reflects how well an FL technique ranks faulty (relevant) methods near the top.
+Precision at a given rank $k$ ($P@k$) is defined as the proportion of faulty methods among the top $k$ methods:
+$P@k=\frac{\text{number of faulty methods in top }k\text{ ranks}}{k}$
+For a given bug, the Average Precision (AP) is computed as the average of the precision values at each rank where a faulty method appears:
+$AP=\frac{1}{M}\sum_{k=1}^{N}P@k\cdot\text{rel}(k)$
+where $M$ is the total number of faulty methods, $N$ is the total number of ranked methods, and $\text{rel}(k)$ is 1 if the method at rank $k$ is faulty, and 0 otherwise.
+Finally, Mean Average Precision (MAP) is obtained by averaging the AP values across all bugs in the dataset:
+$MAP=\frac{1}{Q}\sum_{q=1}^{Q}AP_{q}$
+where $Q$ is the total number of faults. A higher MAP indicates that faulty methods are consistently ranked higher by the localization technique, reflecting better overall performance.
 Mean Reciprocal Rank (MRR).
-MRR considers the position where the first buggy file was ranked and takes the mean from all the bug reports.
-It is computed by the following formula, where K is the set of bug reports and $rank_{i}$ is the position of the first buggy method in the rank.
-$MRR=\frac{1}{K}\sum_{i=1}^{K}\frac{1}{rank_{i}}$
+MRR is a ranking metric that evaluates how quickly the first faulty (relevant) method appears in the ranked list.
+For each fault $q$ (out of $Q$ total faults), let $r_{q}$ denote the rank position of its earliest (highest-ranked) faulty method.
+The reciprocal rank for that fault is then $1/r_{q}$.
+Mean Reciprocal Rank (MRR) is calculated as the mean of these reciprocal ranks across all faults:
+$\mathrm{MRR}=\frac{1}{Q}\sum_{q=1}^{Q}\frac{1}{r_{q}}$
+A higher MRR indicates that, on average, faulty methods appear closer to the top of the ranking, reflecting better fault localization performance.
 Top K.
 The Top K metric represents the number of bugs in which at least one buggy method is located between the first K best-ranked methods in the approach.
-According to a previous study (Parnin and Orso, 2011), developers only check a limited amount of suspicious locations.
+According to a previous study (Parnin and Orso, 2011), developers only check a limited number of suspicious locations.
 Due to this, we use 1, 3, and 5 as the values for K.
-5 Case Study Results
-In this section, we discuss the results of our research questions (RQs). For each RQ, we present the motivation, our approach and the results.
+4 Results and Analysis
+This section discusses the results of our research questions (RQs). We present the motivation, approach, and results for each RQ.
 RQ1 - Are the test failures related to the bug in crash reports?
 Motivation.
-Prior research (Abreu et al., 2007; Jones and Harrold, 2005; Naish et al., 2011) uses SBFL to assist in debugging, which relies on failing test coverage.
-The assumption behind SBFL is that the system has at least one failing test covering the bug (i.e.
-fault-triggering failing test).
-Intuitively, if fault-triggering tests fail on the bug report commits (the nearest commit when a bug report was reported), we can leverage SBFL techniques to pinpoint the buggy locations.
-However, it is not clear if fault-triggering tests are available and failing when a bug report is created (Just et al., 2014a; Haben et al., 2023).
-Therefore, in this RQ, we study whether there are fault-triggering tests (and if these tests fail) at the moment of the bug report creation and their impact on the SBFL’s efficiency.
+Spectrum-Based Fault Localization (SBFL) relies on failing test coverage to identify faulty code, under the assumption that failing tests actually reveal the presence of a bug (Abreu et al., 2007; Jones and Harrold, 2005; Naish et al., 2011).
+However, when a bug report is created, it is often unclear whether the tests that fail at that point are actually related to the bug described in the report-these are known as fault-triggering tests (Just et al., 2014a; Haben et al., 2023; Rafi et al., 2025).
+This question is especially important for crash-related bugs, where developers may only have the stack trace and available test results to guide debugging.
+With this in mind, in RQ1, we investigate whether the test failures observed at the bug report commit are actually caused by the bug described in the crash report.
+We then also explore how the presence or absence of such fault-triggering tests impacts the effectiveness of SBFL in localizing the fault.
+While analyzing fault-triggering tests that are later added to bug reports would reflect post-localization efforts, our goal in RQ1 is to focus on the reporting stage, characterizing the absence of such tests when the bug is first reported, and suggesting a practical solution for this early debugging scenario.
 Approach.
 To understand whether the test failures are related to the bugs in crash reports, we analyze the number of crash reports that contain failing fault-triggering tests.
-As discussed in Section 4.1.2, for each bug report, we perform our study on the bug report commits to avoid biases of ”future” tests.
-In particular, we execute the fault-triggering tests provided by Defects4J and collect the test results with GZoltar.
-We also investigate whether fault-triggering tests are related to other types of bug reports (without stack traces) and compare the test results with crash reports.
-A prior study (Chen et al., 2023) identified a set of bugs without stack traces from Defects4J.
-Therefore, we used their dataset and identified a total of 157 bugs without stack tracks.
-This data will allow us to draw a comparison between crash report bugs and bugs without stack traces.
-To study the impact of missing failing fault-triggering tests on SBFL techniques, we select Ochiai (Abreu et al., 2007) as the baseline since it is widely used and has been shown to perform well on real faults (Pearson et al., 2017; Le et al., 2013; Abreu et al., 2006).
-We use the detailed coverage and the test results obtained via the Gzoltar execution to apply the Ochiai formula at the method level.
-The formula calculates the suspiciousness of each code statement, which allows us to rank the methods based on how probable they contain the bug.
+As discussed in Section 3.1.2, we perform our study on the bug report commits for each bug report to capture the exact state of the project when the crash was first reported and to avoid biases of tests added after the bug was reported.
+Specifically, we execute all available tests using GZoltar to collect test execution results and coverage data, then compare the coverage of failing tests with the methods mentioned in the crash stack traces and the actual buggy methods identified in the fixes.
+Building on the prior work by Rafi et al.
+(2025), which revisited Defects4J to analyze fault-triggering tests and bug timelines, we further inspected the bug reports to determine whether they contain stack traces.
+Through this analysis, we identified 157 bugs without stack traces, allowing us to compare crash-report bugs with non-crash bugs.
+To study the impact of missing fault-triggering tests on SBFL techniques, we select Ochiai (Abreu et al., 2007) as the baseline since it is widely used and has been shown to perform well on real faults (Pearson et al., 2017; Le et al., 2013; Abreu et al., 2006).
+We utilize the detailed coverage and test results obtained through the Gzoltar execution to apply the Ochiai formula at the method level.
+The formula calculates the suspiciousness of each code statement, allowing us to rank the methods based on their likelihood of containing the bug.
 The SBFL suspiciousness formulas vary for each technique, but all are based on the idea that the code that is more covered by failing tests and less covered by passing tests is more likely to be buggy.
@@
 $n_{11}$ yes failed
-in which the terms are as defined in Table  2.
+in which the terms are defined in Table  2.
 Each term n in the formula (e.g., $n_{00}$) corresponds to the number of tests that obey specific criteria for the method j for which the suspiciousness score is being calculated.
@@
   • 6 failing test cases that cover M1 ($n_{11}$)
-  • 2 failing test cases that do not cover M1 ($n_{10}$)
-  • 10 passing test cases that cover M1 ($n_{01}$)
+  • 2 failing test cases that do not cover M1 ($n_{01}$)
+  • 10 passing test cases that cover M1 ($n_{10}$)
 By applying Equation 1, we get that the Ochiai score for method M1 is approximately 0.530.
 Results.
-Only 3.33% of the crash report bugs contain fault-triggering tests (failing test cases that uncover the bug).
+Only 3.33% of the crash report bugs include failing tests that are directly caused by the bug (i.e., fault-triggering tests).
 Table 3 shows the test execution and fault-triggering test results for the crash report bugs.
@@
   • Crash Reports: 2 out of 60 bugs have fault-triggering tests (3.33%)
-  • Bugs without Stack Traces: 16 out of 157 bugs have fault-triggering tests (10.19%)
+  • Bugs without Stack Traces: 16 out of 157 bugs have fault-triggering tests 10.19%)
 Table 3: Test execution results for the crash report bugs.
@@
 Total (60) 24,302 5.987 0.025 2
-Based on the findings, we observed that the majority of the crash report bugs (96.67%) either lack fault-triggering tests or do not trigger them (i.e., fault-triggering tests did not fail).
-In addition, we can see that the percentage of fault-triggering tests in crash report bugs is significantly less when compared to the set of bugs without stack traces (3.33% versus 10.19%).
-The reason may be that such exception-related issues are more often triggered during production and not during testing.
-Another reason may be that developers are more likely to handle exceptions for debugging purposes rather than fault prevention (Shah et al.
-(2008)).
-This shift in the use of exception handling as debugging aids increases the likelihood of exception-related issues occurring in production.
-In such cases, applying traditional SBFL techniques would not be ideal, given the incapacity of the tool to differentiate buggy statements when there are no test failures covering them.
-Due to the absence of failing fault-triggering tests, the runtime information included in the bug report (i.e., stack traces) serves as the final resource for assistance.
-In short, studying how to leverage stack traces for fault localization is an important supplement to traditional SBFL techniques that use failing fault-triggering tests.
-SBFL techniques may not be effective in addressing crash reports due to the limited presence of fault-triggering failing tests (3.33% of all bug reports).
-To study how the absence of fault-triggering tests impacts SBFL, we apply Ochiai to our studied bugs with stack traces.
-To evaluate the approach, we utilize the metrics presented in Section 4.2, specifically Top-1, Top-3, Top-5, MAP and MRR.
-We exclude from the Ochiai results the bugs without failing tests, given that the presence of failing tests is necessary for the computation of the suspicioness score.
+Based on the findings, we observed that most crash report bugs (96.67%) either lack fault-triggering tests or do not trigger them (i.e., fault-triggering tests did not fail).
+In addition, we can see that the percentage of fault-triggering tests in crash report bugs is significantly lower (3.33%) compared to the set of bugs without stack traces (10.19%).
+The reason may be that such exception-related issues are more often triggered during production, rather than during testing.
+Another reason may be that developers are more likely to handle exceptions for debugging purposes rather than fault prevention (Shah et al., 2008).
+This shift in exception handling as a debugging aid increases the likelihood of exception-related issues occurring in production.
+In such cases, applying traditional SBFL techniques would not be ideal, given the tool’s inability to differentiate between buggy and non-buggy statements when no test failures cover them.
+Due to the absence of failing fault-triggering tests, the runtime information included in the bug report (i.e., stack traces) is the final resource for assistance.
+In short, studying how to leverage stack traces for fault localization is an essential supplement to traditional SBFL techniques that use failing fault-triggering tests.
+Only 3.33% of crash-report bugs have failing tests that expose the bug (i.e., fault-triggering tests), limiting the applicability of coverage-based fault localization.
+To study how the absence of fault-triggering tests impacts SBFL, we apply Ochiai to our studied bugs.
+To evaluate the approach, we utilize the metrics presented in Section 3.2, specifically Top-1, Top-3, Top-5, MAP, and MRR.
+We exclude bugs without failing tests from the Ochiai results, given that the presence of failing tests is necessary to compute the suspiciousness score.
 Out of 60 bugs, 23 did not have any failing tests.
-Note that even if there is a failing test, the test may not be related to the bug that we are interested in (i.e., it is not a fault-triggering test).
+Note that even if a test fails, it may not be related to the bug we are interested in (i.e., it is not a fault-triggering test).
 On the bug report commit, SBFL performed poorly in all the projects, only being able to locate 2/60 bugs among the Top-5 methods.
@@
 Ochiai could not locate any bug in Top-1 and located only 1 in Top-3 and 2 in Top-5.
-The highest MAP and MRR values across projects are 0.0693 and 0.0836, respectively, both in the JacksonCore project.
+These two localized bugs are from the JacksonCore and Compress projects, with JacksonCore containing a fault-triggering test.
+The highest MAP and MRR values across projects are 0.0693 and 0.0836, respectively, in the JacksonCore project.
 Considering that the values of MAP and MRR range between 0 and 1, the obtained values are extremely low.
@@
 The MAP and MRR values assess the effectiveness of fault localization techniques in returning relevant results in the top ranking.
-These metrics are important indicators that demonstrate the usefulness of fault localization techniques.
+These metrics are essential indicators that demonstrate the usefulness of fault localization techniques.
 According to a prior survey by Kochhar et al.
 (2016), 80% of developers consider a fault localization technique successful if it can localize bugs in the top 5 positions.
-Hence, our finding shows the inefficacy and limitation of SBFL techniques due to the lack of failing tests related to the bug reports.
+Hence, our finding shows the inefficiency and limitations of SBFL techniques due to the lack of failing tests related to the bug reports.
 Table 4: Ochiai results for all the studied systems.#Bugs represents the total number of studied bugs, while #Bugs with FT is the number of bugs in which there is at least one Failing Test.
@@
 Time (2 — 2) 0 0 0 0.0003 0.0003
-Due to a lack of failing fault-triggering tests, traditional SBFL techniques have inferior localization results.
+Due to a lack of fault-triggering tests, traditional SBFL techniques have inferior localization results.
 RQ2 - What is the relationship between stack traces and buggy location?
 Motivation.
-In the previous RQ, we observed that most crash report bugs do not have fault-triggering tests.
-Because of that, the stack traces stand out as the best source of execution information.
-In this RQ, we want to study how stack traces are related to the buggy location and how they can be used to locate these bugs.
-To do so, we look at two things: (i) what is the type of modification performed in the bug fix to resolve the exception, and (ii) how far away the buggy methods (i.e., methods updated in the bug fix) are from the methods in the stack traces.
+In the previous RQ, we observed that most crash report bugs lack fault-triggering tests.
+As a result, stack traces are the only information available to reflect the code execution at the time of failure.
+In this RQ, we investigate the relationship between the stack traces included in bug reports and the actual buggy code.
+Specifically, we aim to assess whether stack traces provide meaningful guidance for identifying the faulty location and understanding the nature of the fix.
+To do this, we analyze two dimensions: (i) the type of modification made in the fix since this reflects how developers interpreted and responded to the exception, and (ii) the proximity between the modified (buggy) methods and the stack trace methods.
+These dimensions allow us to evaluate the extent to which stack traces directly support fault localization and resolution.
 Approach.
@@
 Results.
-In 83% of the cases, the bugfix intention is to prevent the exception from happening again.
-Figure 7 shows the bugfix intention type distribution for studied bugs.
+In 83% of cases, the bug fix intention is to prevent the exception from happening again.
 In the majority of the fixes (83.3%), the intention behind the fix is Exception Prevention (i.e., to prevent the exception from happening again).
-This means that the stack traces are directly related to the root cause of the bug, which suggests that stack trace is an essential source of information to locate the bug.
+This means that the stack traces are directly related to the root cause of the bug, suggesting that the stack trace is an essential source of information for locating the bug.
 The following two predominant categories are Exception Wrapping (10.0%) and Exception Conversion (5.0%), which also possess a correlation between the fix and the stack traces.
 In only one bug, COMPRESS-31, the exception-related code was not touched for the fix (Keep Throwing category).
-After a more detailed investigation, we noticed that, despite this bug being in the Defect4j database, its resolution field was marked as “Not A Problem” on Jira, which explains why nothing was done about the exception.
-Figure 7: Bugfix intention type distribution in the studied bugs.
+After a more detailed investigation, we noticed that, despite the bug being in the Defect4j database, its resolution field was marked as ‘Not A Problem‘ on Jira, which explains why no action was taken regarding the exception.
 In 78.3% of the bugs, the buggy methods are directly reachable from the stack trace, with an average distance of 0.34 method calls.
@@
 Furthermore, a total of 78.3% of the bugs contain buggy methods that are reachable from the stack trace (i.e., distance is zero or more), averaging a very short distance from the stack trace methods of 0.34 method calls.
-In such scenarios, leveraging stack traces to reconstruct the execution path at the time of the exception can be an effective method for pinpointing the bug.
-We conduct a manual analysis of the remaining 21.7% of the bugs in which the bugs methods are unreachable from the stack trace.
-We find that, in these cases, the stack traces only contain external entries (i.e., entries referring to external libraries) or the provided stack traces were incomplete.
-In short, our findings highlight the potential of using stack traces for fault localization.
-In 98.3% of the studied bugs, the bugfix intention is directly correlated with the exception in the stack trace (Exception Prevention, Exception Conversion or Exception Wrapping).
-In addition, 78.3% of the buggy methods are reachable from the stack traces, having an average distance of 0.34 method calls.
+In such scenarios, the short distance between the stack trace and buggy methods greatly reduces the search space for developers or automated tools.
+This makes stack traces particularly effective for fault localization, as they often point to or near the actual faulty method, even in the absence of explicit failing tests.
+This observation supports our proposed SBEST approach, which uses stack trace methods to approximate fault-triggering behavior.
+We conduct a manual analysis of the remaining 21.7% of the bugs for which the buggy methods are unreachable from the stack trace.
+In these cases, we find that the stack traces only contain external entries (i.e., entries referring to external libraries) or the provided stack traces were incomplete.
+In summary, our findings underscore the potential of utilizing stack traces for fault localization.
+These findings indicate that stack-trace proximity is not only common but also practically useful as a substitute for inadequate test coverage in spectrum-based fault localization when such tests are absent.
+In 98.3% of the studied bugs, the developer’s fix was applied to one of the methods shown in the exception’s stack trace (Exception Prevention, Exception Conversion, or Exception Wrapping).
+Additionally, 78.3% of the buggy methods are reachable from the stack traces, with an average distance of 0.34 method calls.
 This shows that the stack traces are a valuable source of information about the bug.
@@
 Motivation.
-In RQ1, we found that the existing failing tests rarely matched the actual fault-triggering tests.
+In RQ1, we found that the existing failing tests rarely matched the fault-triggering tests.
 Without fault-triggering tests, the performance of SBFL techniques is greatly affected.
@@
 The intention behind the bugfix is usually related to the exception, and the buggy methods are often a short distance away from the stack traces.
-Prior studies (Chen et al., 2021a, b) show that developers usually rely on stack traces when investigating for bugs, as they provide essential information about the buggy location.
+Prior studies (Chen et al., 2021a, b) show that developers usually rely on stack traces when investigating bugs, as they provide essential information about the buggy location.
 Stack traces, similar to fault-triggering tests, carry contextual and execution information associated with the root causes of bugs.
-In a way, stack traces can be used to represent the coverage of a failing fault-triggering test.
+In a way, stack traces can represent the coverage of a failing fault-triggering test.
 Therefore, in this RQ, we aim to investigate how stack traces can be leveraged to complement test cases in locating bugs.
+We conduct this analysis on the same set of 60 crash-reporting bugs identified earlier.
 Approach.
-To understand how stack trace information can contribute to locating buggy methods, we proposed to evaluate an approach called SBEST (Spectrum-Based localization Enhanced by Stack Traces).
-SBEST is a fault localization approach based on SBFL that incorporates stack trace information with the test coverage data.
+To understand how stack trace information can contribute to locating buggy methods, we propose to evaluate an approach called SBEST (Spectrum-Based localization Enhanced by Stack Traces).
+SBEST is a fault localization approach based on SBFL, incorporating stack trace information with the test coverage data.
 This technique applies SBFL principles, but instead of using the failing fault-triggering test coverage in the Ochiai formula, it considers the methods that appear on the stack trace entries as the source of the failure.
-We define a method that is causing a test failure if it appears on any of the frames in a stack trace.
-Prior studies (Wong et al., 2014a; Chen et al., 2021b) have also shown that the position of methods in the stack trace can be useful in fault localization.
-Therefore, we incorporate a Stack Trace (ST) score into our approach.
-We compare our approach with two baselines.
-We use Ochiai as the first one because it is shown to be one of the best SBFL techniques and performs very well on real faults (Pearson et al., 2017).
-In addition, we leverage the position of the stack trace as the second baseline.
-For this, we rank the top entries in the stack trace as more suspicious.
-Below, we describe our approach in detail.
+We define a process causing a test failure if it appears on any of the frames in a stack trace.
+Prior studies (Wong et al., 2014; Chen et al., 2021b) have also shown that the position of methods in the stack trace can be helpful in fault localization.
+Therefore, our approach incorporates a Stack Trace (ST) score.
+We compare our approach with four baselines.
+The first is Ochiai, a widely adopted and high-performing SBFL technique (Pearson et al., 2017).
+The second is the stack trace position baseline, where methods appearing higher in the stack trace are ranked as more suspicious.
+The third is Call Frequency-Based Fault Localization (CFBFL) (Vancsics et al., 2021), which replaces binary hit-based coverage with the frequency of method appearances in unique call stack contexts.
+CFBFL leverages execution information to capture deeper relationships between stack traces and faulty behavior, making it suitable even when explicit failing tests are unavailable or weak.
+To fairly incorporate this method into our evaluation, we first examined the program versions that immediately precede the bug report, consistent with our SBEST setup.
+We then implemented CFBFL using its best-performing variant, JaccardC, as reported in their original study (Vancsics et al., 2021).
+Lastly, we evaluate the SB_score, a partial version of our approach that incorporates coverage from stack-trace-related methods but excludes stack trace position weighting.
 Our overall approach, SBEST, consists of the sum of two scores: the SB score and the ST score.
 Spectrum Based (SB) Score.
-The first one, the SB score, is based on an approach designed to use the information from the stack traces in the fault localization process.
-It is based on the Ochiai formula, and it utilizes existing test coverage information.
-However, instead of using the failing tests as the hint of the bug location (since there may not be any test failure or the failure is not related to the reported bug, as found in RQ1), we use the tests covering the methods in the stack trace as the fault-triggering tests.
-We base this change on the fact that the stack trace represents the system execution at the moment of the failure in the same way that the fault-triggering tests do.
-More specifically, we utilize the methods present on the first five internal entries from the stack traces to select the failing tests that will be applied to Ochiai’s formula.
-We base this design on the fact that the higher the stack trace entry, the more probable it is to be related to the bug (Wong et al., 2014b).
-The selection of failing tests is determined by counting the number of lines each existing test covers in the Stack Trace methods.
+The SB score is based on an approach designed to utilize information from stack traces in the fault localization process.
+It is based on the Ochiai formula and utilizes existing test coverage information.
+However, instead of using the failing tests as a hint of the bug location (since there may not be any test failure or the failure is not related to the reported bug, as found in RQ1), we use the tests covering the methods in the stack trace as the fault-triggering tests.
+While we use Ochiai in this study due to its strong performance and wide adoption, our approach is not tied to this specific formula; it is general and can be applied using any SBFL formula.
+Our key insight is to use the stack trace as a proxy for failing test behavior by identifying tests that cover the methods appearing in the stack trace.
+These tests are then treated as fault-triggering inputs, enabling the application of SBFL even in the absence of actual failing tests.
+While we use the Ochiai formula in our evaluation due to its strong performance and popularity, our approach is not tied to any specific formula.
+The technique is general and can be applied with any SBFL formula, as the novelty lies in how we construct coverage-like information from stack traces rather than in the formula itself.
+We base this change on the fact that the stack trace represents the system’s execution at the moment of failure, in the same way that the fault-triggering tests do.
+More specifically, we utilize the methods present in the first five internal entries from the stack traces to select the proxy failing tests that will be applied to Ochiai’s formula.
+We base this design on the fact that the higher the stack trace entry, the more probable it is to be related to the bug (Wong et al., 2014).
+The selection of proxy failing tests is determined by counting the number of lines each existing test covers in the Stack Trace methods.
 We first compute the number of covered lines from these methods for each test by applying the following formula:
 $ST\_Covered\_lines\_number_{t}=\sum_{m=1}^{5}CL_{m}$ (2)
-in which m represents each of the top 5 methods in the Stack Trace, and $CL_{m}$ is the number of lines in the correspondent method that is covered by the test.
-Then, we select the set of the X highest results by applying the following formula, in which T represents the set of tests and X is a threshold for the number of selected tests:
-$\begin{split}T_{\text{failing}}=\{t\in T\mid t\text{ is one of the }X\text{ tests with the highest }\\ ST\_Covered\_lines\_number_{t}\}\end{split}$ (3)
-We noticed that selecting a value for X that was too low made the approach often miss buggy locations because it was too focused on just a few methods.
-On the other hand, a high threshold also generates inaccurate results since the approach is unable to differentiate between the methods in the stack trace.
-Upon experimentation, we noticed that setting it to 15 gave the best results.
+in which m represents each of the top 5 methods in the Stack Trace, and $CL_{m}$ is the number of lines in the corresponding method covered by the test.
+Then, we select the set of X highest results by applying the following formula, in which T represents the set of tests and X is a threshold for the number of selected tests:
+[…]
+We noticed that selecting a value for X that was too low often missed buggy locations because it was too focused on just a few methods.
+On the other hand, a high threshold also generates inaccurate results since the approach cannot differentiate between the methods in the stack trace.
+Experimenting showed that setting it to 15 gave the best results.
 Finally, we apply the Ochiai formula (1) using these 15 tests as failing tests and all the remaining as passing tests.
+Table 6 presents the experimental results for different values of proxy failing tests X, which will be discussed further in the results section.
 Stack Trace (ST) Score.
 The SB score described above merges the information in the stack traces with the test coverage to compute the suspiciousness score for a given method.
-In addition to that, we make use of the ST score used in a previous study (Chen et al., 2021b) in order to boost the stack trace’s impact into the calculated score.
-To do this, after having computed the SB_score following the approach described before, we sum to it a ST_score calculated in the following manner:
+In addition, we use the ST score used in a previous study (Chen et al., 2021b) to boost the stack trace’s impact on the calculated score.
+To do this, after having computed the SB_score following the approach described before, we add to it a ST_score calculated in the following manner:
 $ST\_score=\begin{cases}\frac{1}{ST\_rank}&\text{if }ST\_rank\leq 10\\ 0.1&\text{if }ST\_rank>10\\ 0&\text{if }\text{method not found}\end{cases}$ (4)
-In which ST_rank is the position in which the given method appears in the stack trace after the external entries are removed.
+In which ST_rank is the position where the given method appears in the stack trace after the external entries are removed.
 For example, if the method appears in the second position, the ST_score would be 1/2=0.5.
-Suspiciouness Score. After having both the STCB Ochiai score and the ST Score, we compute the suspiciousness score of a giving method by applying the formula:
-$Suspiciouness\_score=STCB\_Ochiai\_score+ST\_score$ (5)
-Suspiciouness Rank.
+Suspiciousness Score. After having both the SB Ochiai score and the ST Score, we compute the suspiciousness score of a given method by applying the formula:
+Suspiciousness_score=SB_Ochiai_score+ST_scoreSuspiciousness\_score=SB\_Ochiai\_score+ST\_score (5)
+Suspiciousness Rank.
 After having the final suspiciousness score, we generate the final suspiciousness rankings.
 The higher the suspiciousness score, the higher the ranking.
-For instance, if the highest calculated suspiciousness score is 1.0 for the method M8, M8 is going to be the first method in the suspiciousness rank, as it has the highest probability of containing the bug.
+For instance, if the highest calculated suspiciousness score is 1.0 for the method M8, M8 will be the first method in the suspiciousness rank, as it has the highest probability of containing the bug.
 Table 5: Fault localization results for all the studied projects. #Bugs represents the total number of studied bugs in each project.
    System (#Bugs)    Technique    Top-1    Top-3    Top-5    MAP    MRR
-      Cli (2)       Ochiai       0       0       0       0.0045       0.0045
+   Cli (2)    Ochiai    0    0    0    0.005    0.005
    Stack Trace    1    1    2    0.6    0.6
-      SB_score       0       0       1       0.1333       0.1333
+   SB_score    0    0    1    0.133    0.133
+   CFBFL    0    0    1    0.222    0.222
    SBEST    1    1    2    0.6    0.6
-      Closure (8)       Ochiai       0       0       0       0.0007       0.0012
-      Stack Trace       1       2       2       0.1710       0.2170
+   Closure (8)    Ochiai    0    0    0    0.001    0.001
+   Stack Trace    1    2    2    0.171    0.217
    SB_score    0    0    0    0.0003    0.005
-      SBEST       1       2       2       0.1462       0.1671
+   CFBFL    0    0    1    0.006    0.0056
+   SBEST    1    2    2    0.146    0.167
    Codec (1)    Ochiai    0    0    0    -    -
-      Stack Trace       1       1       1       0.3333       1
-      SB_score       1       1       1       0.3893       1
-      SBEST       1       1       1       0.5853       1
-      Compress (11)       Ochiai       0       0       1       0.0079       0.023
-      Stack Trace       3       6       8       0.4122       0.4425
-      SB_score       0       3       3       0.1404       0.1587
-      SBEST       3       9       9       0.4620       0.4904
+   Stack Trace    1    1    1    0.333    1
+   SB_score    1    1    1    0.389    1
+   CFBFL    1    1    1    0.399    1
+   SBEST    1    1    1    0.585    1
+   Compress (11)    Ochiai    0    0    1    0.008    0.023
+   Stack Trace    3    6    8    0.412    0.443
+   SB_score    0    3    3    0.141    0.159
+   CFBFL    0    1    1    0.002    0.002
+   SBEST    3    9    9    0.462    0.491
    Csv (1)    Ochiai    0    0    0    -    -
-      Stack Trace       0       1       1       0.3333       0.3333
+   Stack Trace    0    1    1    0.333    0.333
    SB_score    1    1    1    1    1
+   CFBFL    0    1    1    0.014    0.019
    SBEST    1    1    1    1    1
    Gson (3)    Ochiai    0    0    0    -    -
-      Stack Trace       1       2       2       0.5196       0.5196
-      SB_score       1       1       1       0.3338       0.3338
-      SBEST       1       2       2       0.5101       0.5101
-      JacksonCore (4)       Ochiai       0       1       1       0.0693       0.0836
+   Stack Trace    1    2    2    0.52    0.52
+   SB_score    1    1    1    0.334    0.334
+   CFBFL    0    0    0    0.0001    0.0001
+   SBEST    1    2    2    0.51    0.51
+   JacksonCore (4)    Ochiai    0    1    1    0.069    0.084
    Stack Trace    1    1    1    0.25    0.25
-      SB_score       0       0       0       0.0409       0.0457
-      SBEST       0       1       1       0.1609       0.1634
-      JacksonDatabind (10)       Ochiai       0       0       0       0.0005       0.0005
-      Stack Trace       1       2       3       0.1686       0.2446
-      SB_score       0       1       1       0.0125       0.0212
-      SBEST       1       3       3       0.0359       0.0777
+   SB_score    0    0    0    0.041    0.046
+   CFBFL    0    0    1    0.001    0.001
+   SBEST    0    1    1    0.161    0.164
+   JacksonDatabind (10)    Ochiai    0    0    0    0.001    0.001
+   Stack Trace    1    2    3    0.169    0.245
+   SB_score    0    1    1    0.013    0.021
+   CFBFL    1    1    1    0.176    0.215
+   SBEST    1    3    3    0.036    0.078
    Jsoup (8)    Ochiai    0    0    0    0    0
-      Stack Trace       3       3       5       0.4410       0.4444
-      SB_score       0       0       0       0.0214       0.0318
-      SBEST       3       4       4       0.4070       0.4279
+   Stack Trace    3    3    5    0.441    0.445
+   SB_score    0    0    0    0.022    0.032
+   CFBFL    2    3    4    0.224    0.299
+   SBEST    3    4    4    0.407    0.428
    JxPath (1)    Ochiai    0    0    0    -    -
-      Stack Trace       1       1       1       0.3333       0.3333
-      SB_score       0       0       0       0.0625       0.0625
+   Stack Trace    1    1    1    0.333    0.333
+   SB_score    0    0    0    0.063    0.063
+   CFBFL    0    0    0    0.001    0.001
    SBEST    1    1    1    1    1
-      Lang (4)       Ochiai       0       0       0       0.0008       0.0008
-      Stack Trace       1       3       3       0.4375       0.4375
-      SB_score       0       1       1       0.1411       0.1411
-      SBEST       2       3       3       0.5835       0.5835
+   Lang (4)    Ochiai    0    0    0    0.001    0.001
+   Stack Trace    1    3    3    0.438    0.438
+   SB_score    0    1    1    0.141    0.141
+   CFBFL    0    0    1    0.01    0.01
+   SBEST    2    3    3    0.584    0.584
    Math (3)    Ochiai    0    0    0    0.0002    0.0001
-      Stack Trace       2       2       2       0.2215       0.4583
-      SB_score       0       0       0       0.0428       0.0292
-      SBEST       2       2       2       0.1372       0.3472
-      Mockito (2)       Ochiai       0       0       0       0.0060       0.0281
-      Stack Trace       0       1       1       0.0238       0.1667
-      SB_score       0       1       1       0.0371       0.25
-      SBEST       0       1       1       0.0371       0.25
+   Stack Trace    2    2    2    0.222    0.458
+   SB_score    0    0    0    0.041    0.029
+   CFBFL    0    1    1    0.113    0.129
+   SBEST    2    2    2    0.137    0.347
+   Mockito (2)    Ochiai    0    0    0    0.006    0.028
+   Stack Trace    0    1    1    0.024    0.167
+   SB_score    0    1    1    0.037    0.25
+   CFBFL    0    0    0    0.002    0.002
+   SBEST    0    1    1    0.037    0.25
    Time (2)    Ochiai    0    0    0    0.0003    0.0003
-      Stack Trace       0       1       2       0.2917       0.2917
-      SB_score       1       1       1       0.5001       0.5001
-      SBEST       0       1       1       0.3333       0.3333
-      Total (60)       Ochiai       0       1       2       0.00902       0.01421
-      Stack Trace       16       27       34       0.32406       0.40992
-      SB_score       4       10       11       0.20396       0.26517
-      SBEST       17       32       33       0.42846       0.49647
+   Stack Trace    0    1    2    0.292    0.292
+   SB_score    1    1    1    0.5    0.5
+   CFBFL    0    0    0    0.001    0.001
+   SBEST    0    1    1    0.333    0.333
+   Total (60)    Ochiai    0    1    2    0.009    0.014
+   Stack Trace    16    27    34    0.324    0.41
+   SB_score    4    10    11    0.204    0.265
+   CFBFL    4    8    13    0.084    0.136
+   SBEST    17    32    33    0.429    0.497
+Table 6: Fault Localization Performance with Different Numbers of Failing Tests for SBEST
+   X (Proxy Failing Tests)    Top-1    Top-3    Top-5    MAP    MRR
+   5    14    22    27    0.37111    0.43887
+   10    16    29    30    0.39121    0.43944
+   15    17    32    33    0.42846    0.49647
+   20    14    23    28    0.37532    0.42598
+   25    12    20    25    0.30897    0.41720
+Table 7: Fault Localization Performance for Different Numbers of Top Stack-Trace Methods for SBEST
+Top #Stack-trace Methods Top-1 Top-3 Top-5 MAP MRR
+5 17 32 33 0.42846 0.49647
+10 7 10 11 0.38962 0.41266
+15 4 7 9 0.20011 0.2409
 Results.
@@
 Table 5 presents the metrics results for the 2 baselines as well as for the SB_score alone and the SBEST approach for all the systems.
-We can see that the worst-performing technique is Ochiai, which, as discussed before, locates only 2 bugs in Top-5.
+We can see that the worst-performing technique is Ochiai, which, as discussed before, locates only 2 bugs in the Top 5.
 The best-performing approach is SBEST, but the Stack Traces ranking is a close second.
-The Stack Trace ranking locates 16 bugs in Top-1, 27 in Top-3 and 34 in Top-5.
+The Stack Trace ranking locates 16 bugs in Top-1, 27 in Top-3, and 34 in Top-5.
 Surprisingly, the Stack Traces alone provide very good results, locating more than half of the bugs (56.67%) within the Top 5.
-This shows that the stack traces are a very important source of information about the bug location and should be prioritized when it comes to studying FL in crash reports.
-The SB_score is able to locate some bugs that the Stack Trace ranking does not locate, but in general, it does not perform very well, identifying only 18.33% of the buggy methods in the Top 5.
-We can see, however, that the SB_score is substantially improved by the addition of the ST_score.
-The resultant technique SBEST performs slightly better than the Stack Traces in the Top 1 and Top 3, locating 17 and 32 bugs, respectively, and locates only one bug less in the Top 5, totalling 33 bugs spotted within the Top 5 rank.
-This technique and the Stack Traces ranking take turns as the best localization approach for each project.
-For example, the Stack Trace ranking is better overall for the JacksonCore project, while SBEST has the top performance on the Lang bugs.
-In total, SBEST locates 17 bugs in Top-1, 32 in Top-3 and 33 in Top-5.
-The average MAP and MRR for SBEST are 0.42846 and 0.49647, respectively, representing a 32.22% improvement on the MAP and a 17.43% in the MRR when compared to the Stack Traces.
-These results show that the stack traces can complement existing passing test information to help locate the bug root causes.
-Overall, out of 60 studied bugs, SBEST successfully located 17 bugs in Top-1, 32 in Top-3 and 33 in Top-5.
-In addition, it achieves an improvement of 32.22% on MAP and 17.43% on MRR when compared to the Stack Trace ranking.
-6 Discussion and Implication of Our Findings
+This shows that stack traces are a significant source of information about bug location and should be prioritized when studying FL in crash reports.
+The SB_score can locate some bugs that the Stack Trace ranking does not locate, but in general, it does not perform very well, identifying only 18.33% of the buggy methods in the Top 5.
+CFBFL, which ranks methods based on their frequency in unique stack contexts during execution, identified 4 bugs in Top-1, 8 in Top-3, and 13 in Top-5.
+While it improves over Ochiai, it lags behind Stack Trace ranking and SBEST in all metrics.
+The MAP and MRR of CFBFL (0.084 and 0.136) are significantly lower than those of SBEST and Stack Traces, suggesting that while call frequency provides useful contextual signals, it may not be sufficient on its own for localizing crash faults when fault-triggering test coverage is sparse or missing.
+We observe that adding the ST score improves the SB_score.
+The resulting technique, SBEST, outperforms the Stack Traces approach in the Top 1 and Top 3 ranks, locating 17 and 32 bugs (compared to Stack Traces’ 16 and 27), and only one fewer bug in the Top 5, for a total of 33.
+Overall, SBEST and Stack Traces are the two most effective localization techniques across all projects.
+Stack Traces performs better on JacksonCore, while SBEST achieves the best results on Lang.
+On average, SBEST attains a MAP of 0.42846 and an MRR of 0.49647, showing moderate improvements (32.22% and 17.43%) over the Stack Trace baseline in scenarios where fault-triggering tests are unavailable.
+These results indicate that stack traces can effectively complement passing test information in identifying root causes.
+Overall, out of the 60 studied bugs, SBEST successfully located 17 bugs in the Top 1, 32 in the Top 3, and 33 in the Top 5.
+Additionally, it achieves improvements of 32.22% on MAP and 17.43% on MRR compared to the Stack-Trace-based ranking.
+SBEST performs best with 15 proxy failing tests and the top 5 stack trace methods.
+Tables 6 and 7 present the fault localization performance of SBEST under different parameter settings for the number of proxy failing tests and the number of stack trace methods, respectively.
+FL performance is stable between 10-20 proxy failing tests, with Top-1 results peaking at 17 faults.
+However, the performance degrades when we include more than 15 proxy failing tests or more than 5 stack trace methods.
+This suggests that, when too many failing tests or stack-trace methods are included, more noise may be introduced, which can negatively impact fault localization accuracy.
+For example, using 25 proxy failing tests or 15 stack methods significantly reduces Top-1 and MAP scores.
+Setting the number of proxy failing tests (X) to 15 and using the top 5 stack trace methods gives the best results, with 17 bugs in Top-1 and a MAP of 0.4285.
+Other configurations show a drop in performance.
+5 Discussion and Implications of Our Findings
 On the use of stack traces in the absence of failing tests.
-From the results obtained in Section 5, we observe that the majority of the crash report bugs do not contain fault-triggering tests.
-In such cases, the test failure information is not available for the fault localization process, causing techniques such as SBFL to fail.
-The stack traces, on the other hand, demonstrate to be deeply correlated to the bug cause.
+From the results obtained in Section 4, we observe that most crash report bugs do not contain fault-triggering tests.
+In such cases, the test failure information is unavailable for the fault localization process, causing techniques such as SBFL to fail.
+Conversely, the stack traces are deeply correlated with the cause of the bug.
 We found that, in 66% of the bugs, at least one buggy method is directly listed in the stack trace.
 Even when this is not the case, the buggy methods are usually just a few calls away from it.
-In addition, the most common bugfix intention is to prevent the exception in the stack trace from happening again, which highlights the big association between them.
+In addition, the most common bugfix intention is to prevent the exception in the stack trace from happening again, highlighting the significant association between them.
 Leveraging stack trace rankings for enhanced fault localization.
 In the real world, test failures are not always present.
-Especially when it comes to production-phase bugs, it is important to look for alternative sources of information for fault localization.
-Making use of the data available in logs and stack traces shows to be a good path to take in such scenarios.
-Even though previous studies have used stack traces for Fault Localization, many are based on information retrieval (IRFL) approaches (Zhou et al., 2012; Saha et al., 2013; Lam et al., 2017).
-IRFL techniques treat all the content in the bug report, including the stack traces, as textual information, therefore missing important context such as the ranking of the stack traces.
-Our results show that the stack traces ranking alone was able to locate more than half of the bugs within Top-5 and, thus, should be better utilized.
-Encouraging the integration of stack traces to complement execution information.
+Especially regarding production-phase bugs, it is essential to look for alternative sources of information for fault localization.
+Using the data available in logs and stack traces is a good path for such scenarios.
+Although previous studies have utilized stack traces for Fault Localization, many are based on information retrieval (IRFL) approaches (Zhou et al., 2012; Saha et al., 2013; Lam et al., 2017).
+IRFL techniques treat all content in the bug report, including stack traces, as textual information; therefore, they miss important context, such as the ranking of the stack traces.
+Our results show that the stack traces ranking alone could locate more than half of the bugs within Top-5 and, thus, should be better utilized.
+Encouraging the use of stack traces to complement execution information.
 SBEST, our approach combining the stack trace ranking with the code coverage, shows promising results.
-The improvement in the metrics implies that the coverage information was able to enhance the stack trace ranking, aiding in obtaining a better overview of the system execution in the moment of failure.
-Our main goal when studying this technique was to gain a deeper understanding of the scenario of these bugs and how each piece of information available can help with fault localization.
-We believe that this is just the first step and that future research can benefit from the aforementioned findings to build more intricate techniques.
+The improvement in the metrics implies that the coverage information enhanced the stack trace ranking, aiding in obtaining a better overview of the system execution at the moment of failure.
+Our primary goal in studying this technique was to gain a deeper understanding of the scenario surrounding these bugs and how each piece of available information can aid in fault localization.
+We believe this is just the first step, and that future research can benefit from the findings above to develop more sophisticated techniques.
 In addition, we make all the data from this study available at Zenodo [^1] https://zenodo.org/records/11062413, including the detailed code coverage and test results of the bug report commits that took us 100 hours to obtain.
-We think that this data will be useful for future research on subjects such as fault localization, automatic bug fixing, test generation, etc.
-Our data does not contain post-bug-fixing development knowledge, which will allow researchers to tailor tools more aligned with the real-world scenario.
+We believe this data will be beneficial for future research on fault localization, automatic bug fixing, test generation, and other related areas.
+Our data excludes any post-fix developer knowledge, which would allow researchers to tailor tools more closely aligned with real-world scenarios.
+Limitations.
+While our findings highlight the potential of stack traces for fault localization in the absence of fault-triggering tests, we acknowledge several limitations.
+Our study focuses specifically on bugs that result in exceptions with available stack traces.
+In practice, many bugs do not manifest as crashes or produce exceptions, which may limit the applicability of our findings to the broader space of software faults.
+Future work should investigate the applicability of stack-trace-based localization to a wider range of bug types and software systems.
+Future Directions.
+Our findings open several avenues for future research.
+One particularly important direction is to explore how execution signals such as stack traces and runtime logs can be better leveraged in the absence of fault-triggering tests.
+These lightweight signals can serve as a proxy for fault-triggering execution, allowing fault localization methods like SBFL to operate in real-world scenarios where fault-triggering tests are missing.
+Additionally, future studies could expand evaluation beyond Defects4J to improve generalizability across projects and programming languages.
+Some recent research, such as AutoFL (Kang et al., 2024b), applies LLMs to reason over code and coverage data for fault localization.
+At the same time, LIBRO (Kang et al., 2024a) generates bug-reproducing tests directly from natural language bug reports with stack traces.
+Building on these ideas, one direction is to leverage LLMs to interpret the underlying causes of failures from stack traces and coverage data, and use that information to filter out irrelevant methods or code regions.
+This could enable more accurate and efficient ranking of suspicious methods, potentially supporting automated patch generation and thereby enhancing the practicality of SBEST in scenarios where explicit test failures are absent.
 6 Threats to Validity
@@
 6.1 External Validity.
-Threats to external validity relate to how generalizable our findings are.
-To minimize this threat, we conduct our studies in 15 systems from Defects4j, a famous framework of real bugs used in multiple other studies.
-These systems are widely used and vary in size and number of tests.
-Even though every system is written in Java, our method is not restricted to Java-based systems.
-In addition, despite the low percentage of bugs with stack traces in Defects4j, this aligns with the percentage reported in other systems (Chen et al., 2021a).
-Moreover, we opted not to use the buggy version made available in defects4j but the code available at the moment of the bug report creation in order to better emulate the scenario in which the developers investigate the bug.
-Finally, even though we opted for applying only Ochiai when analyzing SBFL’s efficiency when applied to crash report bugs, it is shown to be one of the best SBFL techniques and performs very well on real faults (Pearson et al., 2017).
+Threats to external validity relate to the generalizability of our findings.
+To minimize this threat, we selected 15 systems from Defects4J, as used by  Just et al.
+(2014b), a widely used benchmark of real bugs in prior studies.
+These systems vary in size, test coverage, and domain.
+However, we acknowledge that our study focuses on 60 crash-report bugs ($\approx$10% of Defects4J) with available stack traces.
+This proportion aligns with observations from real-world studies, which show that only a limited subset of user-reported bugs contain stack traces (Bettenburg et al., 2008; Chen et al., 2021a).
+Although this represents a subset of all bugs, crash reports are among the most practically significant cases developers face, as stack traces often provide the first diagnostic evidence when failures occur and have been widely utilized in fault localization research (Bettenburg et al., 2008; Zou et al., 2021).
+We prioritized crash-reporting bugs for which stack traces and test coverage could be reliably extracted, though this was sometimes challenging due to legacy and dependency issues.
+We publicly release our dataset to support replication and further study.
+Finally, while our evaluation focuses on Java projects, the proposed approach can be extended to other programming languages in future work.
 6.2 Internal validity.
@@
 In this study, we only analyze Defects4J bugs.
-Although Defects4j is widely used (Lutellier et al., 2020; Chen et al., 2021b; Li et al., 2021; Ye et al., 2022), the findings might be different in other systems, specially non-Java systems.
-Another threat to internal validity is the way in which we selected the bug reports with stack traces.
+Although Defects4J is widely used (Lutellier et al., 2020; Chen et al., 2021b; Li et al., 2021; Ye et al., 2022), the findings might be different in other systems, especially non-Java systems.
+Another threat to internal validity is how we selected the bug reports with stack traces.
 We collect stack traces in the description and comments, which is the typical place in which they are added.
@@
 Spectrum-based fault Localization (SBFL) approaches are highly recognized due to their accuracy, efficiency, and simplicity.
-However, their effectiveness depends on the availability of fault-triggering failing tests.
-In the lack of such test failures, alternative approaches should be considered.
-In this paper, we research using stack traces for SBFL in the absence of failing tests.
+However, their effectiveness depends on the availability of fault-triggering tests.
+In the absence of such test failures, alternative approaches should be considered.
+In this paper, we investigate the use of stack traces for SBFL in the absence of fault-triggering tests.
 We study 60 crash reports from the Defects4J benchmark.
-We find that most of the studied crash reports do not contain fault-triggering failing tests.
+We find that most of the studied crash reports do not contain fault-triggering tests.
 This results in very low efficiency when using traditional SBFL.
@@
 In addition, in 78.3% of the bugs, the buggy method is directly reachable from the stack trace, with an average distance of only 0.34 method calls.
-This shows that the buggy methods are usually very close to the stack traces when it comes to the execution call graph.
-On top of that, our results show that even without any advanced technique, the stack traces alone provide a good indication of the buggy locations, being capable of locating more than half of the bugs on Top-5.
-Finally, we develop a simplified SBFL method called SBEST that uses the stack trace information in place of the failing tests and thus integrates it into the coverage information to perform the SBFL.
-SBEST was able to locate 17 bugs in Top-1, 32 in Top-3 and 33 in Top-5, which represents an improvement of 32.22% on MAP and 17.43% on MRR when compared to the Stack Trace ranking.
-The enhancement of the metrics suggests that the coverage information improved the stack trace ranking and helped the technique gain a clearer picture of the system’s execution at the time of failure.
-We believe that future studies can benefit from the findings of this research to develop more intricate techniques.
+This indicates that buggy methods are typically very close to the stack traces in terms of the execution call graph.
+Furthermore, our results demonstrate that even without employing advanced techniques, the stack traces alone provide a reliable indication of buggy locations, capable of locating more than half of the bugs in the Top 5.
+Finally, we develop a simplified SBFL method called SBEST that utilizes stack trace information as a proxy for failing tests, thereby integrating it with coverage information to perform SBFL.
+SBEST located 17 bugs in Top-1, 32 in Top-3, and 33 in Top-5, with MAP and MRR values that are 32.22% and 17.43% higher than the Stack Trace ranking, respectively.
+The results suggest that combining coverage information with stack traces can enhance fault localization by offering a clearer view of system execution when fault-triggering tests are unavailable.
+Future studies can benefit from the findings of this research to develop more intricate techniques.
 In addition, all the data that we have made available can help studies in areas such as automatic bug fixing, test generation, and, of course, fault localization.
-As potential future work, we plan to expand our analysis by increasing our bugs dataset utilizing other sources beyond Defects4j.
+As potential future work, we plan to expand our analysis by increasing our bug dataset by utilizing other sources beyond Defects4J.
+8 Declarations
+Conflicts of Interest.
+All authors are from Concordia University and the University of Alberta. There may be conflicts with individuals from these two institutions.
+Funding.
+The authors have no funding to report.
+Data Availability Statements.
+We make the data publicly available online for replication and encourage future research in this area (Rafi et al., 2024).